### v2

# Revisiting Differentially Private Hyper-parameter Tuning

Abstract
We study the application of differential privacy in hyper-parameter tuning, a crucial process in machine learning involving selecting the best hyper-parameter from several candidates. Unlike many private learning algorithms, including the prevalent DP-SGD, the privacy implications of tuning remain insufficiently understood or often totally ignored. Recent works propose a generic private selection solution for the tuning process, yet a fundamental question persists: is this privacy bound tight?
This paper provides an in-depth examination of this question. Initially, we provide studies affirming the current privacy analysis for private selection is indeed tight in general. However, when we specifically study the hyper-parameter tuning problem in a white-box setting, such tightness no longer holds. This is first demonstrated by applying privacy audit on the tuning process. Our findings underscore a substantial gap between current theoretical privacy bound and the empirical bound derived even under strong audit setups.
This gap motivates our subsequent investigations. Our further study provides improved privacy results for private hyper-parameter tuning due to its distinct properties. Our results demonstrate broader applicability compared to prior analyses, which are limited to specific parameter configurations.
Index Terms:
Hyper-parameter Tuning; Privacy Audit; Private Selection; Differential Privacy;
1 Introduction
Differential Privacy (DP) [18] stands as the prevailing standard for ensuring privacy in contemporary machine learning. A ubiquitous technique employed to ensure DP across a diverse array of machine learning tasks is differentially private stochastic gradient descent (DP-SGD, a.k.a., noisy-SGD) [6, 45, 2]. The ongoing refinement of DP-SGD’s privacy analysis [2, 4, 33, 35], referred to as “privacy accounting” or “privacy budgeting”, has contributed to a profound comprehension of the mechanism.
In addition to a single (private) training process, machine learning systems always involve a hyper-parameter tuning process that entails running a (private) base algorithm (e.g., DP-SGD) multiple times with different configurations and selecting the best run. Regrettably, unlike the well-studied DP-SGD, the reasoning for the privacy cost of such tuning operations is inadequately studied and often totally ignored.
Naively, one can bound the privacy loss for the tuning operation by the composition theorem. If we run the private base algorithm $k$ times with different hyper-parameters, the total privacy cost deteriorates at most linearly with $k$ (or $\mathcal{O}(\sqrt{k})$ if the base algorithm satisfies approximate DP and we use advanced composition theorem [20], which is nearly optimal [24]). However, these bounds are still far from satisfactory as $k$ is usually large in practice. Perhaps due to this limitation, it still remains common to exhaustively tune a private algorithm to achieve strong performance but only consider the privacy cost for a single run [15, 56, 54, 43].
For another approach, tuning hyper-parameter privately can be framed as a private selection problem, for which several well-explored mechanisms, such as the sparse vector technique [19] and the exponential mechanism [32], may potentially be utilized. However, these mechanisms assume that the score function (defining the “best” to be selected) has low sensitivity (for DP analysis), which is a condition not always met.
Thanks to the contributions of Liu and Talwar [30], hyper-parameter tuning now enjoys significantly better privacy bound than naively applying composition theorem. To briefly describe their findings, if we run a private base algorithm a random number of times (possibly with different hyper-parameters) and only output the best single run, the privacy cost only deteriorates by a constant multiplicative factor [30]. For example, if the base algorithm is $(\varepsilon,0)$-DP, then the whole tuning process is $(3\varepsilon,0)$-DP if the running number follows a geometric distribution [30]. This is much better than the $(k\varepsilon,0)$-DP bound under fixed $k$ times of running.
Later, Papernot and Steinke [42] operate within Rényi DP (RDP) framework [33] and mandate the randomization of the number of running times, presenting additional results for varying degrees of randomness. A noteworthy aspect of both methodologies [30, 42] lies in their treatment of the base algorithm as a black box; thus, such a generic approach applies to a broader spectrum of private selection problems, provided the base algorithm is differentially private on its own.
Motivations. Although hyper-parameters can be tuned with rigorous privacy, whether the enhanced privacy analysis [30, 42] is tight or not is still an open problem. We note that results provided by [30] and [42] for certain setups show that the privacy cost still increases substantially (e.g., tripling compared to the base algorithm’s privacy cost as shown above); however, it also seems plausible that only revealing the best single run should not consume that much privacy budget. This prompts a natural question: Does hyper-parameter tuning truly consume a noticeably larger privacy budget than the base algorithm, as indicated, for instance, by a factor of roughly three [30]?
A negative answer would meaningfully imply that a more nuanced privacy-utility trade-off can be achieved through analysis alone. Conversely, a positive answer would indicate significant improvement in analysis is impossible. Overall, in contrast to the well-established understanding of privacy deterioration due to composition, investigating such a problem contributes to a deeper understanding of how privacy degrades due to an alternative factor: selection.
This work. We answer the posed question with both positive and negative answers. In the affirmative, our constructed example demonstrates that the current generic privacy bound provided in [42] for private selection is indeed tight. Still, the result only holds in the worst case. Conversely, in the negative, we uncover a more favorable privacy bound given the base algorithm is DP-SGD specifically. Our contributions are as follows.
1) Validating tightness of generic privacy bound for private selection (Section 3.2). We first provide a private selection instance where we observe only a negligible gap between the true privacy cost (with our detailed analysis) and the cost predicted by the current privacy bound [42]. However, when we study the private hyper-parameter tuning problem, where the base algorithm is DP-SGD, such tightness no longer holds. This finding is related to our other two contributions.
2) Empirical investigation on the privacy leakage of hyper-parameter tuning (Section 4). We first take empirical approaches to investigate how much privacy is leaked when performing hyper-parameter tuning. This is done via the privacy audit technique [39, 24, 23], an interactive protocol used to empirically measure the privacy of some mechanisms. In contrast, unlike all previous privacy auditing work, which focuses on the privacy of the base algorithm (e.g., DP-SGD), auditing the tuning procedure is a fresh problem that requires new formulation and insight. Specifically, the score function, used to select the “best”, is the new factor that needs to be settled.
We formulate various privacy threat models tailored for hyper-parameter tuning, where the weakest one corresponds to the most practical scenario and the strongest one corresponds to the worst case. Results under the weakest provide evidence that the tuning process hardly incurs additional privacy costs beyond the base algorithm. Notably, even the empirical privacy bound derived from the strongest adversary still exhibits a substantial gap from the generic privacy bound proposed by [42]. This gap motivates us to derive better theoretical results in the remaining sections.
3) Improved privacy results (Sections 5.2, 5.3 and 6). We find that tuning a DP-SGD protocol does enjoy a better privacy result, as demonstrated by our subsequent study on deriving an improved privacy result. The pivotal aspect driving this improvement lies in representing the privacy of the base algorithm with finer resolution, and DP-SGD does have a distinctive characterization. This is done within the $f$-DP framework [17], deviating from the well-known $(\varepsilon,\delta)$-DP [18] or RDP [33]. We show that our results are tight in a general sense. Our results are also more generalizable, contrasting to previous work [30, 42], where they only easily apply to a limited range of parameter setups.
Subsequent to our improved results is a further experimental evaluation. We aim to compare our improved privacy result with the empirical privacy lower bound derived under an idealized audit setup. Notably, there is still a gap in between. This finding is examined in detail, revealing that the score function, a new factor in auditing hyper-parameter tuning, is a key determinant influencing audit performance. To give a concise summary of our main contributions:
  • We confirm the general tightness of the current generic privacy bound for private selection.
  • We empirically evaluate the privacy leakage due to hyper-parameter tuning.
  • We provide improved privacy results for private hyper-parameter tuning.
2 Background
2.1 Differential Privacy (DP)
Definition 1 (Differential Privacy [18]).
Given a data universe $\mathcal{X}$, two datasets $X,X^{\prime}\subseteq\mathcal{X}$ are adjacent if they differ by one data example. A randomized algorithm $\mathcal{M}$ satisfies $(\varepsilon,\delta)$-differential privacy, or $(\varepsilon,\delta)$-DP, if for all adjacent datasets $X,\,X^{\prime}$ and for all events $S$ in the output space of $\mathcal{M}$, we have $\operatorname{Pr}(\mathcal{M}(X)\in S)\leq e^{\varepsilon}\operatorname{Pr}(\mathcal{M}(X^{\prime})\in S)+\delta$.
Differentially private algorithms are resilient to post-processing, and the execution of multiple DP algorithms sequentially, known as composition, also maintains DP. Rényi DP (RDP), a DP relaxation shown in the following, often serves as a tight analytical tool to assess the privacy cost under composition.
Definition 2 (Rényi DP [34]).
The Rényi divergence is defined as $\mathcal{D}_{\alpha}(M||N)=\frac{1}{\alpha-1}\ln\mathbb{E}_{x\sim N}\left[\frac{M(x)}{N(x)}\right]^{\alpha}$ with $\alpha>1$. A randomized mechanism $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$ is said to be $(\alpha,\gamma)$-Rényi DP, or $(\alpha,\gamma)$-RDP, if $\mathcal{D}_{\alpha}(\mathcal{M}(X)||\mathcal{M}(X^{\prime}))\leq\gamma$ holds for any adjacent dataset $X,X^{\prime}$.
Differentially private stochastic gradient descent (DP-SGD) [6, 45, 2]. A machine learning model denoted as $f_{w}$ often represents a neural network with trainable parameters $w$. Specifically tailored for classification tasks in this study, $f_{w}$ may take image data as input and output corresponding labels. Stochastic Gradient Descent (SGD) [28] is the method for updating $w$ iteratively. Hyper-parameters, such as the updating step size (a.k.a. learning rate), often need to be tuned to gain optimized performance.
As the private version of SGD, DP-SGD ❶ computes the per-sample gradient for each sub-sample data, ❷ clips each gradient to have bounded $l_{2}$ norm, and ❸ adds Gaussian noise (Gaussian mechanism). The resultant private gradient $p_{i}$ is used to update parameter $w$, and $i$-th (up to $N$-th) update can be expressed as:
$p_{i}=\sum_{(x,y)\in\mathbf{B}}\operatorname{CLP}_{C}\left(\nabla_{w}\ell\left(w_{i-1};x,y\right)\right)+R_{i}$ (1)
$w_{i}\leftarrow w_{i-1}-\mathbf{lr}\cdot p_{i}$
where $\mathbf{lr}$ is the learning rate and data batch $\mathbf{B}$ contains the sub-sampled data (e.g., $x$ may represent an image, and $y$ is its label) where the sample ratio is $\tau$. The function $\operatorname{CLP}_{C}(u)=u\cdot\min(1,\frac{C}{\|u\|_{2}})$ where the clipping threshold $C$ is a hyper-parameter. $R_{i}$ is calibrated isotropic Gaussian noise sampled from $\mathcal{N}(0,C^{2}\sigma^{2}\mathbb{I}^{d})$ where $d$ is the dimension or number of trainable parameters and $\sigma$ is the noise multiplier. $\ell(;)$ is the loss function that represents the neural network $f_{w}$ and some loss metric (e.g. cross-entropy loss). With $R_{i}$ and $\operatorname{CLP}_{C}$ operation eliminated, we recover SGD. There are notable variants of DP-SGD, such as DP-Adam [48]; they share the same privacy analysis as DP-SGD because of the post-processing property of DP.
2.2 Privacy Auditing
Hypothesis testing interpretation of DP. For a randomized mechanism $\mathcal{M}$, let $X,X^{\prime}$ be adjacent datasets, let $y\in\mathcal{Y}$ be the output of $\mathcal{M}$ taking input $X$ or $X^{\prime}$, we form the null and alternative hypotheses:
$\mathbf{H_{0}}:\text{$X$ was the input},\text{\quad}\mathbf{H_{1}}:\text{$X^{\prime}$ was the input}.$ (2)
For any decision rule $\mathcal{R}:\mathcal{Y}\rightarrow\{0,1\}$ in such a hypothesis testing problem, it has two notable types of errors: 1) type I error or false positive rate $\mathrm{FP}=\operatorname{Pr}(\mathcal{R}(y)=1|\mathbf{H_{0}})$, i.e., the probability of rejecting $\mathbf{H_{0}}$ while $\mathbf{H_{0}}$ is true; 2) type II error or false negative rate $\mathrm{FN}=\operatorname{Pr}(\mathcal{R}(y)=0|\mathbf{H_{1}})$, i.e., the probability of rejecting $\mathbf{H_{1}}$ while $\mathbf{H_{1}}$ is true. DP can be characterized by such two error rates as follows.
Theorem 1 (DP as Hypothesis Testing [24]).
For any $\varepsilon>0$ and $\delta\in[0,1]$, a mechanism $\mathcal{M}$ is $(\varepsilon,\delta)$-DP if and only if
$\mathrm{FP}+\mathrm{e}^{\varepsilon}\mathrm{FN}\geq 1-\delta,\text{\quad}\mathrm{FN}+\mathrm{e}^{\varepsilon}\mathrm{FP}\geq 1-\delta$ (3)
both hold for any adjacent dataset $X,X^{\prime}$ and any decision rule $\mathcal{R}$ in a hypothesis testing problem as defined in Equation (2).
Theorem 1 has the following implications. With $\delta$ fixed at some value, under the threat model that an adversary can only operate at some $\mathrm{FP}$ and $\mathrm{FN}$ under some decision rule $\mathcal{R}$ for a specific adjacent dataset pair $X,X^{\prime}$, a lower bound
$\varepsilon_{L}^{(X,X^{\prime},\mathcal{R})}=\max\{\log\frac{1-\delta-\mathrm{FP}}{\mathrm{FN}},\log\frac{1-\delta-\mathrm{FN}}{\mathrm{FP}},0\}$ (4)
can be computed, meaning that the algorithm cannot be more private than that, i.e., the true privacy parameter $\varepsilon_{T}\geq\varepsilon_{L}^{(X,X^{\prime},\mathcal{R})}$, just as entailed by Theorem 1. Finding $\varepsilon_{T}$ requires taking the maximum of lower bound value over all pairs of $X,X^{\prime}$ and $\mathcal{R}$, which is clearly intractable in general. In practice, people are satisfied by reporting an upper bound $\varepsilon_{U}\geq\varepsilon_{T}$, which is obtained by analytical approaches (privacy accounting) [2, 33, 35].
Algorithm 1 Game-based Privacy Audit $\mathcal{G}$
1:DP protocol $\mathcal{P}$, adjacent pair $X,X^{\prime}$
2:$b_{\text{truth}}\leftarrow\{0,1\}$triangleright Trainer flips a fair coin
3:$\hat{X}\leftarrow X$ if $b_{\text{truth}}=0$, $\hat{X}\leftarrow X^{\prime}$ otherwise
4:Run $\mathcal{P}(\hat{X})$triangleright Trainer runs the private protocol
5:$b_{\text{guess}}\leftarrow\{0,1\}$triangleright Adversary makes a guess based on $\mathcal{P}(\hat{X})$
6:$(b_{\text{truth}},b_{\text{guess}})$
Privacy audit. Privacy audit aims to find a lower bound of the privacy cost for a private protocol $\mathcal{P}$ based on the hypothesis testing interpretation of DP as shown above. This is usually done via simulating the interactive game-based protocol described in Algorithm 1. Such a simulation is typically repeated many times, resulting in many pairs of $(b_{\text{truth}},b_{\text{guess}})$. Then, the $\mathrm{FP}$ and $\mathrm{FN}$ for adversary’s guessing are computed by Clopper-Pearson method [11], which means that lower bound $\varepsilon_{L}^{(X,X^{\prime},\mathcal{R})}$ computed by Equation (4) is with a confidence specification. If the adversary can make very accurate guesses and derive a lower bound higher than some claimed privacy parameter, it suggests $\mathcal{P}$ is not private as claimed.
Related work on privacy audit. In privacy-preserving machine learning, privacy audit mainly serves a different goal from that of certain earlier studies [52, 16, 8, 7] on detecting privacy violation in general query-answering applications. Previous work on privacy audit in machine learning mainly targets auditing the DP-SGD protocol to assess its theoretical versus practical privacy [39, 24, 23]. Additional studies [47, 37, 31, 57] concentrate on enhancing the strength of audits on DP-SGD (yielding stronger/larger-value lower bound) or improving the efficiency (incurring fewer simulation overheads). Drawing a parallel to the action of guessing whether a data point was included or not, privacy audit may also be linked to membership inference attack (MIA) [38, 44], but diverges from MIA in terms of both their goals and focus. There are also recent works on auditing prediction [9] and synthetic data generation [3], which differ from our auditing experiments.
2.3 Private Hyper-parameter Tuning
Problem Formulation. We formulate the private hyperparameter tuning problem aligning with [30, 42]. Let $\Omega=\{\mathcal{M}_{1},\mathcal{M}_{2},\cdots,\mathcal{M}_{m}\}$ be a collection of DP-SGD algorithms. These correspond to $m$ possible hyper-parameter configurations. We have $\mathcal{M}_{i}:\mathcal{X}\rightarrow\mathcal{Y}$ for $i\in[m]$, and all of these algorithms satisfy the same privacy parameter, i.e., they are all $(\varepsilon,\delta)$-DP for the same $\varepsilon,\delta$. The practitioner can freely determine the size of $\Omega$.
The goal is to return an algorithm element (including its execution) of $\Omega$ such that the output of such algorithm has (approximately) the maximum score as specified by some score function $g:\mathcal{Y}\rightarrow\mathbb{R}$. The score function $g$ usually serves a utility purpose (e.g., $g$ could evaluate the validation loss on a held-out dataset). The selection must be performed in a differentially private manner. The general private selection problem corresponds to the cases where $\Omega$ contains some arbitrary differentially private algorithms.
Related work on private hyper-parameter tuning. Well-established algorithms like the sparse vector technique [19] and exponential mechanism [32] may potentially be leveraged to the tuning problem; however, they assume a low sensitivity in the metric defining the “best”, a condition not always applicable. Some earlier work [10] also suffers from the same issue. Papernot et al. [42] and Liu [30] have provided generic private selection approaches circumventing such challenges. Mohapatra et al. [36] study privacy issues in adaptive hyper-parameter tuning under DP, which is different from the non-adaptive tuning problem considered in this work. There is related work [25] that builds upon [42]; therefore, what we understand about the generic approach in this study also applies to [25].
Focus of this paper. Our first focus is to leverage and formulate specific privacy audit instantiation to understand how privacy deteriorates due to selection, diverging from all previous privacy audit work on privacy deteriorating due to composition. Furthermore, we also study improving privacy results specifically for a white-box application: the hyper-parameter tuning problem, pre-conditioned on the base algorithm is DP-SGD.
Algorithm 2 Private Selection Protocol $\mathcal{H}$
1:Dataset $X$; algorithms $\Omega$; distribution $\xi$; score function $g$
2:Draw a sample: $k\leftarrow\xi$
3:$Y\leftarrow\text{{Null}}$, $S\leftarrow-\infty$
4:for $i=1,2,\cdots,k$ do
5:    Uniformly randomly fetch one element $\mathcal{M}_{i}$ from $\Omega$
6:    $y_{i}\leftarrow\mathcal{M}_{i}(X)$triangleright Run $\mathcal{M}_{i}$ on dateset $X$
7:    If $g(y_{i})>S$: $Y\leftarrow y_{i}$, $S\leftarrow g(y_{i})$triangleright Selecting the “best”
8:end for
9:$Y$
3 Current Private Selection Protocol
3.1 The Private Selection Algorithm
The state-of-the-art algorithm [30, 42] for private selection is outlined in Algorithm 2. Notably, this generic algorithm can be applied as long as the base algorithm is differentially private on its own. When each element $\mathcal{M}_{i}$ inside $\Omega$ is a DP-SGD instance, we have our private hyper-parameter problem.
Current privacy analysis. Specifically, if the base DP algorithm $\mathcal{M}$ is $(\varepsilon,0)$-DP (pure DP) and $\xi$ is a geometric distribution, Algorithm 2 is $(3\varepsilon,0)$-DP [30]. An improved bound for the pure DP case is provided in [42] by replacing the distribution $\xi$ with the Truncated Negative Binomial (TNB) distribution at some specific parameter setups, as shown in Appendix A.2. This improvement is achieved through RDP analysis.
3.2 Our General Tightness Proof
We show the current privacy bound due to [42] is tight in a general sense. Our contribution of providing the following example is that the tightness we prove is non-asymptotic, unlike the tight example shown in [42], which relies on assumptions and approximations.
Example 1 (Our Construction for Pure DP).
Let $\mathcal{M}$ have a finite output space $\mathcal{Y}=\{\mathrm{A},\mathrm{B},\mathrm{C}\}$. $\mathcal{M}$ only cares about the number of data samples in its input. If the number is even, its output follows the distribution shown as the left-hand side of Equation (5); otherwise, its output distribution is the right-hand side.
$\operatorname{Pr}_{\mathcal{M}}\left\{\begin{array}[]{ll}\operatorname{Pr}_{\mathrm{A}}=1-b\mathrm{e}^{\varepsilon}-db\\ \operatorname{Pr}_{\mathrm{B}}=b\mathrm{e}^{\varepsilon}\\ \operatorname{Pr}_{\mathrm{C}}=db\\ \end{array}\right.\operatorname{Pr}_{\mathcal{M^{\prime}}}\left\{\begin{array}[]{ll}\operatorname{Pr}_{\mathrm{A}^{\prime}}=1-b-db\mathrm{e}^{\varepsilon}\\ \operatorname{Pr}_{\mathrm{B}^{\prime}}=b\\ \operatorname{Pr}_{\mathrm{C}^{\prime}}=db\mathrm{e}^{\varepsilon}\\ \end{array}\right.$ (5)
where $\operatorname{Pr}_{\mathrm{A}}$ denotes the probability of event $A$ occurs conditioned on even (similarly we also have $\operatorname{Pr}_{\mathrm{A}^{\prime}}$ with respect to odd). With $b=10^{-3},d=10^{2},\varepsilon=1$, we can see $\mathcal{M}$ is clearly $(1,0)$-DP for any pair of adjacent (w.r.t. addition/removal) dataset.
Let each element $\mathcal{M}_{i}$ fetched from $\Omega$ in line 5 of Algorithm 2 has the same output distribution as Equation (5). Also let a score function $g$ give $g(\mathrm{C})>g(\mathrm{B})>g(\mathrm{A})$. Let $\xi$ be the TNB distribution with parameter $\eta=1,\nu=10^{-3}$ (geometric distribution). The probability for each event that Algorithm 2 outputs is computed by the following.
Claim 1.
Let $y$ be some event in $\mathcal{Y}$, the probability of $y$ occurs as the output of the tuning process $\mathcal{H}$ (Algorithm 2) is
$\operatorname{Pr}(y)=\sum_{k\sim\xi}\operatorname{Pr}(k)\left(\operatorname{Pr}(E_{\leq y})^{k}-\operatorname{Pr}(E_{<y})^{k}\right),$ (6)
where $E_{\leq y}=\{x:g(x)\leq g(Y)\}$ and $E_{<y}=\{x:g(x)<g(Y)\}$. See proof in Appendix B.1.
Let $\operatorname{Pr}_{\mathcal{H}},\operatorname{Pr}_{\mathcal{H^{\prime}}}$ denote the probabilities for each event conditioned on $\mathcal{H}$ operates on adjacent dataset pair. For $\operatorname{Pr}_{\mathcal{H}}$ we have
$\operatorname{Pr}_{\mathcal{H}}\left\{\begin{array}[]{ll}\operatorname{Pr}_{\mathrm{A}|\mathcal{H}}=\sum_{k\sim\xi}\operatorname{Pr}(k)\operatorname{Pr}_{\mathrm{A}}^{k}\\ \operatorname{Pr}_{\mathrm{B}|\mathcal{H}}=\sum_{k\sim\xi}\operatorname{Pr}(k)((\operatorname{Pr}_{\mathrm{A}}+\operatorname{Pr}_{\mathrm{B}})^{k}-\operatorname{Pr}_{\mathrm{A}}^{k})\\ \operatorname{Pr}_{\mathrm{C}|\mathcal{H}}=\sum_{k\sim\xi}\operatorname{Pr}(k)(1-(\operatorname{Pr}_{\mathrm{A}}+\operatorname{Pr}_{\mathrm{B}})^{k})\\ \end{array}\right.$
where $\operatorname{Pr}_{\mathrm{A}|\mathcal{H}}$ denotes the probability of event $A$ occurs as the output of $\mathcal{H}$ conditioned on the input dataset contains even number of data points. $\operatorname{Pr}_{\mathcal{H^{\prime}}}$ can be computed similarly. Numerically, this gives the probabilities shown below
$\operatorname{Pr}_{\mathcal{H}}\left\{\begin{array}[]{ll}\operatorname{Pr}_{\mathrm{A}|\mathcal{H}}=8.66\times 10^{-3}\\ \operatorname{Pr}_{\mathrm{B}|\mathcal{H}}=2.60\times 10^{-4}\\ \operatorname{Pr}_{\mathrm{C}|\mathcal{H}}=9.91\times 10^{-1}\\ \end{array}\right.\operatorname{Pr}_{\mathcal{H^{\prime}}}\left\{\begin{array}[]{ll}\operatorname{Pr}_{\mathrm{A}|\mathcal{H}^{\prime}}=2.66\times 10^{-3}\\ \operatorname{Pr}_{\mathrm{B}|\mathcal{H}^{\prime}}=1.34\times 10^{-5}\\ \operatorname{Pr}_{\mathrm{C}|\mathcal{H}^{\prime}}=9.97\times 10^{-1}\\ \end{array}\right.$ (7)
and it can be checked to satisfy $(2.96,0)$-DP. The theoretical bound claims Algorithm 2 is $(3,0)$-DP, i.e., it is tight up to a negligible gap.
Tightness can also be verified under various $(\eta,\nu)$ setups of TNB. We can also confirm the tightness for approximate DP ($\delta>0$) trivially, as shown in Appendix A.2.
Now a new question arises: Does this tightness shown in the above worst-case still hold for private hyper-parameter tuning where the selection is among several executions of DP-SGD protocol? We investigate this problem in the remaining sections where we hold elements $\mathcal{M}_{i}$ inside $\Omega$ are DP-SGD instances satisfying the same privacy parameter.
Notation Meaning
$\mathcal{G}$ The distinguishing game, Algorithm 1
$\mathcal{P}$ A general protocol to be audited in $\mathcal{G}$
$\mathcal{H}$ The private tuning protocol, Algorithm 2
$\mathcal{M}$ The base algorithm (DP-SGD) of $\mathcal{H}$
$\mathbb{F},\mathbb{M},\mathbb{C},\mathbb{S}$ Datasets used, shown in Section 4.3
$N$ Number of iterations inside $\mathcal{M}$
$C$ Clipping threshold in Equation (1)
$w_{i}$ Model at $i$-th iteration in Equation (1)
$\ell$ The loss function in Equation (1)
$\xi$ Running number distribution of $\mathcal{H}$
$g$ Score function evaluating $\mathcal{M}$’s output
$\mathbf{z}$ Differing data point, constructed by adversary
$p_{i}^{\mathbf{z}}$ $\mathbf{z}$’s gradient at $i$-th iteration in Equation (1)
$p_{i}$ Private gradient in Equation (1)
$\mathbf{Z}_{D}$ Hypothetical $\mathbf{z}$ leading to Dirac gradient
$\lambda_{a},\lambda_{b}$ Two proxies constructed by the adversary
$\sigma$ Noise s.t.d. for $R_{i}$ in Equation (1)
$\varepsilon_{B}$ Base algorithm $\mathcal{M}$’s privacy budget
$\varepsilon_{L}$ Lower bound for $\mathcal{H}$ by audit
$\varepsilon_{U}$ Generic upper bound for $\mathcal{H}$, by [42]
TABLE I: Notations used in our empirical study.
4 Empirical Investigation
In this section, we aim to find how much privacy is leaked due to the tuning procedure $\mathcal{H}$ when the base algorithm is specifically the DP-SGD protocol. Notations used are summarised in Table I.
4.1 High-level Procedure
First, simulate $\mathcal{G}$. We instantiate Algorithm 1 for our experiments, shown in Figure 1. Each execution of $\mathcal{P}$ in $\mathcal{G}$ is an execution of our tuning protocol $\mathcal{H}(\hat{X},\Omega,\xi,g)$. $\Omega$ contains many base algorithms (DP-SGD instances with different hyper-parameter setups) satisfying the same privacy parameter. $\xi$ is the TNB distribution [42] shown in Appendix A.2. $g$ is the score function.
Second, conclude the lower bound. Our null and alternative hypothesis are
$\mathbf{H_{0}}:\text{$X$ was used},\text{\quad}\mathbf{H_{1}}:\text{$X^{\prime}$ was used}.$ (8)
After many simulations of $\mathcal{G}$ where each one gives an assertion for the above hypothesis testing problem, the $\mathrm{FP}$ and $\mathrm{FN}$ are computed by the Clopper-Pearson method [11] with a $95\%$ confidence. We then leverage methods proposed in [37] to compute the empirical privacy lower bound $\varepsilon_{L}^{(X,X^{\prime},\mathcal{R})}$. We provide the detailed procedure for deriving $\varepsilon_{L}^{(X,X^{\prime},\mathcal{R})}$ in Appendix A.3. We omit the notation $(X,X^{\prime},\mathcal{R})$ under clear context.
4.2 Audit Scenario Formulation
This section is to elucidate the four “arrows” originating from the adversary shown in Figure 1.
Forming $X,X^{\prime}$. W.o.l.g., we assume $X^{\prime}=X\cup\{\mathbf{z}\}$. Note that the adversary can always set $X$ to some available datasets. $\mathbf{z}$, known as “canaries” [37], is instantiated as follows.
  • Weaker version. The adversary can select $\mathbf{z}$ to be any real-world data, and to have higher distinguishing performance, $\mathbf{z}$ is set to be sampled from a distribution different from those in $X$.
  • Stonger version. The adversary can directly control the gradient of $\mathbf{z}$, a.k.a., gradient canary. Specifically, it is assumed that adversary generates $\mathbf{z}=\mathbf{Z}_{D}$ such that its gradient is a Dirac vector $\nabla_{w}\ell(w;\mathbf{Z}_{D})=[C,0,0,\cdots,0]^{T}$ [37], i.e., only the first coordinate equals to the clipping threshold $C$ and the rest are all zeros.
Figure 1: Diagram of the distinguishing game $\mathcal{G}$.
Score function $g$. W.o.l.g., the best model is selected if it has the highest score. This new factor distinguishes auditing $\mathcal{H}$ from all previous auditing tasks. We formalize two types of adversaries that are only possible.
  • Weaker version. $g$ is not manipulated, e.g., $g$ is a normal routine to evaluate the model’s accuracy/loss on an untampered validation dataset.
  • Stronger version. The adversary can arbitrarily control $g$, e.g., $g$ can be a routine to evaluate the model’s performance on some malicious dataset.
Adversary’s observation $Y$. Under the assumption of DP-SGD protocol, the whole training trajectory $\{p_{i}\}_{i=1}^{N}$ is released. Equivalently, all the checkpoints $\{w_{i}\}_{i=1}^{N}$ of the neural network are trivially derivable as each checkpoint is just post-processing of the private gradient. Hence, we can denote the observation as $Y=\{p_{1},p_{2},\cdots\,p_{N},w_{1},\cdots,w_{N}\}$. This information corresponds to line 4 of Algorithm 1 or the output of $\mathcal{H}$. Note that including $w_{i},i\in\{1,2,\cdots,N\}$ in $Y$ may be redundant; however, it is for notation convenience as we will later refer to the $w_{i}$ information contained in $Y$.
Figure 2: NTNV setup. Each row corresponds to different types of differing data $\mathbf{z}$; each column corresponds to different training datasets. The vertical axis shows the values for $\varepsilon_{U}$ and audited $\varepsilon_{L}$ based on different proxies.
Adversary’s assertion. Adversary’s assertion is exactly the action shown in line 5 in Algorithm 1. This requires the adversary to transform observations $Y$ into binary guesses. The adversary forms a real-number proxy and compares it to some threshold to make assertions. Proxies formulation will be described in the forthcoming sub-experiments.
Based on the above considerations, we form the following scenarios with increasing levels of threat.
  • Normal training and normal validation (NTNV). The name says that the training dataset is some natural, normal dataset, and the validation for the trained model is also normal, i.e., score function $g$ is not manipulated.
  • Normal training and controlled validation (NTCV). The training dataset is the same as that of NTNV; however, the validation for the trained model is controlled, i.e., score function $g$ is manipulated.
  • Empty training and controlled validation (ETCV). The training dataset is empty (malicious), $g$ is the same as that of NTCV.
4.3 Evaluation Methods
Given the complexity of this subject, here we describe how to evaluate our experimental result. The used dataset and details for running the experiment are provided in Appendix A.3.
For notation convenience, we use abbreviations for the used datasets: $\mathbb{F}$ stands for the FASHION dataset, $\mathbb{M}$ for MNIST, $\mathbb{C}$ for CIFAR10 and $\mathbb{S}$ for SVHN. We use “[]” to fetch the information from some data container. For instance, we use $v[0]$ to denote fetching the first coordinate of $v$ if $v$ is a vector. We also abuse the notation and use $Y[w_{N}]$ to denote fetching the parameter $w_{N}$ from output/observation $Y$.
Results indexing. Our main audit results are presented in figures, and we index them in the following form:
$X=\mathbb{F},\mathbf{z}=\mathbb{M}[I],g=\text{{SF}}[1],$
which means that such a result corresponds to 1) setting $X$ to be the FASHION dataset; 2) setting the differing data $\mathbf{z}$ to be the $I$-th data sample from MNIST dataset; 3) setting the score function $g$ to be the first candidate shown in Table II. Note that $X^{\prime}=X\cup\{\mathbf{z}\}$ and we always shuffle the dataset initially.
Not manipulated 1: $g(Y)=-\sum_{i}\ell(Y[w_{N}];\mathbb{V}[i])$ $\mathbb{V}$ is original validation dataset
Manipulated 2: $g(Y)=-\ell(Y[w_{N}];\mathbf{z})$ 3: $g(Y)=(Y[w_{0}]-Y[w_{N}])[0]$
TABLE II: Score functions are indexed by $\text{{SF}}[a],a=1,2,3$.
Evaluation method. Our main focus is to compare the following bounds; hence, understanding their intuitive interpretations is beneficial.
  • $\varepsilon_{L}$ is the amount of information leakage the adversary can extract based on the execution of $\mathcal{H}$.
  • $\varepsilon_{B}$ is the maximal information leakage due to a single run of the base algorithm, as guaranteed by theoretical analysis [35, 2].
  • $\varepsilon_{U}$ is the maximal information leakage due to execution of $\mathcal{H}$, as guaranteed by theoretical analysis [30, 42].
These bounds are all based on fixed $\sigma$ values. Specifically, after $\sigma$ is fixed for the base algorithm $\mathcal{M}$, we 1) compute $\varepsilon_{B}$ by previous privacy analysis for DP-SGD such as TensorFlow privacy [1]; 2) compute $\varepsilon_{U}$ for $\mathcal{H}$ by current generic bound for hyper-parameter tuning [42]; 3) apply privacy audit to $\mathcal{H}$, obtaining $\varepsilon_{L}$ as shown in Section 4.1. We know that $\varepsilon_{B}\leq\varepsilon_{U}$ is always true, and it is interesting to make the following comparison.
  • $\varepsilon_{L}$ V.S. $\varepsilon_{U}$. This is the main focus. The question to be answered in this comparison is: does hyper-parameter tuning $\mathcal{H}$ practically leak sensitive information ($\varepsilon_{L}$) as predicted by the current generic bound [42] ($\varepsilon_{U}$)?
  • $\varepsilon_{L}$ V.S. $\varepsilon_{B}$. This is another interesting comparison. The question to be answered in this comparison is: How does running a DP-SGD many times and then returning the best (an execution of $\mathcal{H}$) practically leak information ($\varepsilon_{L}$) compared to a single run of DP-SGD ($\varepsilon_{B}$)?
Figure 3: NTCV setup. Sub-figure arrangement is identical to Figure 2.
4.4 Experiments When $g$ Not Manipulated
NTNV, most practical. This scenario corresponds to the most practical setup in our experiments. Experimental results are shown in Figure 2, notated according to Section 4.3. Here the score function is not manipulated.
Assertion. The selection behaves normally, i.e., the best model is selected if it has the highest score (lowest loss) on the original validation dataset. A base proxy $\lambda_{a}$ will be formed following previous work [39, 37] as follows. Compute $\mathbf{z}$’s gradient at each iteration before model update:
$\lambda_{a}=\frac{1}{N}\sum_{i=1}^{N}\frac{1}{C^{2}}\langle p_{i}^{\mathbf{z}},p_{i}\rangle.$ (9)
where $\langle a,b\rangle$ is the inner product of two vectors. By design, $p_{i}^{\mathbf{z}}$ is possibly orthogonal to other independent data’s gradient. Moreover, the adversary has access to the score function. Hence, it seems reasonable to leverage such additional information. To this end, we will also form another proxy $\lambda_{b}$ based on the score function as follows.
$\lambda_{b}=\lambda_{a}-g(Y)$ (10)
$\lambda_{b}$ is our newly formed proxy and can be seen as the enhanced version of $\lambda_{a}$ in auditing hyper-parameter tuning because it tries to include additional information from the score function $g$.
By design, higher value of $\lambda_{a}$ or $\lambda_{b}$ incentivizes the adversary to accept $\mathbf{H_{1}}$. The rationale behind these setups is to expect the abnormally differing data (if $X^{\prime}$ is used, or $\mathbf{z}$ was included in the training) to have a detrimental impact on the training so that the model has a higher loss (lower value of $g(Y)$), making it more distinguishable if $X^{\prime}$ is used ($\mathbf{z}$ was included in the training).
Results. Experimental results are presented in Figure 2, where we present audited $\varepsilon_{L}$ results corresponding to proxy $\lambda_{a}$ or $\lambda_{b}$. We also present the theoretical upper bound $\varepsilon_{U}$ for comparison. An obvious phenomenon is that the audited $\varepsilon_{L}<\varepsilon_{B}<\varepsilon_{U}$ across all setups shown in the first row of Figure 2. The interesting phenomenon is that $\varepsilon_{L}<\varepsilon_{B}$ and the gaps between them are obvious. This means that the adversary cannot even extract more sensitive information than the base algorithm’s (a single run of DP-SGD) privacy budget allows. This shows the adversary’s power is heavily limited under the most practical setting.
In contrast, in Figure 2(e) and Figure 2(f), when the base algorithm’s privacy budget $\varepsilon_{B}=1$, we see that 1) $\varepsilon_{L}$ is greater than the counterpart in the first row of Figure 2, and 2) $\varepsilon_{L}$ is much closer to $\varepsilon_{B}$. This confirms that the differing data $\mathbf{z}$ that has Dirac gradient gains the adversary more distinguishing power than some natural data. We can also observe that $\lambda_{b}$ has almost no advantage over $\lambda_{a}$, indicating additional information from the score provides limited help, at least when the score function $g$ is not manipulated. Another phenomenon is that the audited $\varepsilon_{L}$ under $\varepsilon_{B}=1$ in Figure 2(d) is weaker than that in Figure 2(e) and Figure 2(f), this suggests that auditing performance depends on $X$.
4.5 Experiments When $g$ Manipulated
NTCV, middle-level attack. This scenario corresponds to some middle-level adversary’s power. Experimental results are shown in Figure 3, notated according to Section 4.3. The score function is manipulated, different from that in NTNV.
Assertion. By design, the rationale behind manipulating the score function to be SF[2] is to expect the training to memorize the different data $\mathbf{z}$, and the best model is selected based on this metric. Manipulating the score function to be SF[3] builds on the fact that for $\mathbf{Z}_{D}$, it suffices only to investigate the first coordinate of the model to recover any trace of $\mathbf{z}=\mathbf{Z}_{D}$. The proxy $\lambda_{a}$ is identical to that in Equation (9), however, $\lambda_{b}=\lambda_{a}+g(Y)$ is set in NTCV, which is different from that in NTNV. This is because $g$ is manipulated. Under the same design considerations, a higher value of $\lambda_{a}$ or $\lambda_{b}$ incentivizes the adversary to accept $\mathbf{H_{1}}$.
Results. Experimental results are presented in Figure 3, organized similarly to Figure 2. We observe a phenomenon similar to NTNV that $\varepsilon_{L}$ sees a big gap to $\varepsilon_{B}$ shown in the first row of Figure 3. In contrast, for the results seen in the second row of Figure 3(e) and Figure 3(f), when base algorithm’s privacy budget $\varepsilon_{B}=1$, we have $\varepsilon_{L}\approx\varepsilon_{B}$. Again, this confirms the Dirac gradient canary is more powerful.
As $g$ is manipulated in this case, it is interesting to compare the performance due to $\lambda_{a}$ and $\lambda_{b}$. We can see that $\lambda_{a}$ and $\lambda_{b}$ have almost the same performance, similar to NTNV where $g$ is not manipulated. It gives more evidence that the score itself provides limited additional help.
Figure 4: ETCV setup. Sub-figure arrangement is identical to Figure 2.
Figure 5: ETCV setup when $\mathbf{z}=\mathbf{Z}_{D}$.
ETCV, worst-case. Experimental results are shown in Figure 4 and Figure 5, notated according to Section 4.3. This scenario corresponds to the greatest adversary’s power in our settings. The training dataset is set to be empty, and the score function is manipulated. Assertion. By design, the rationale behind the empty dataset setup is to eliminate the uncertainties due to normal training data’s gradient so that audit performance is maximized, as the adversary only cares about the causal effect from $\mathbf{z}$ to the output [49]. $\lambda_{a},\lambda_{b}$ are set identically to NTCV. Again, higher value of $\lambda_{a}$ or $\lambda_{b}$ incentivize the adversary to accept $\mathbf{H_{1}}$.
Results. Experimental results are presented in Figure 4 and Figure 5. In Figure 4(a), we can see that $\varepsilon_{L}\approx\varepsilon_{B}$ under all setups; we also notice that $\varepsilon_{L}$ still sees a small gap to $\varepsilon_{B}$ under some setups; however, $\varepsilon_{L}$ gets much closer to $\varepsilon_{B}$ compared with that in NTNV and NTCV. The increased audit performance is due to $X=\emptyset$, which eliminates unwanted disturbances for the adversary.
In Figure 5, when $\mathbf{z}=\mathbf{z}_{D}$ is the Dirac gradient canary instead of some natural data, we observe $\varepsilon_{L}\geq\varepsilon_{B}$ under all setups. This suggests that, operationally, hyper-parameter tuning does leak additional privacy beyond what’s allowed to be disclosed by the base algorithm. This also means that tuning hyper-parameters while only accounting the privacy cost for a single run (i.e., naively taking $\varepsilon_{U}=\varepsilon_{B}$ ) is problematic in a rigorous manner. On the other hand, like the results in the previous two setups, we also observe that 1) $\lambda_{a}$ and $\lambda_{b}$ have almost the same performance, and 2) there is a big gap between $\varepsilon_{L}$ and $\varepsilon_{U}$.
4.6 Discussion
Beyond our experimental findings, several noteworthy phenomena emerge, prompting further discussion on the theoretical V.S. practical privacy of hyper-parameter tuning.
Is it safe to tune hyper-parameter while only considering a single run’s privacy cost? The weakest audit setup, representative of practical scenarios, reveals that the tuning procedure hardly leaks more privacy beyond the base algorithm. This suggests a potential safety in tuning hyper-parameters while only accounting for the privacy loss for a single run, aligning with conventional (although not rigorous) practices predating the advent of the generic private selection approach [30, 42]. While our strongest audit reveals noticeable privacy leakage beyond the base algorithm, it is crucial to note that these findings are confined to the stronger adversary, which should be reasonably impractical.
What training dataset benefits the adversary? Comparing results from NTNV/NTCV with results from ETCV, to reach stronger audit result (higher $\varepsilon_{L}$ value), the adversary clearly favors the worst-case setting, i.e., $X=\emptyset,X^{\prime}=\mathbf{z}$ and $\mathbf{z}$ is adversary-chosen. This is because only the causal effect from differing data $\mathbf{z}$ to the output is informative to the adversary [49], and any other independent factors affecting the output will only confuse the adversary to more likely fail the game. Let us re-investigate Equation (1). Specifically, the clipped gradients of data examples other than $\mathbf{z}$ itself introduce additional unwanted uncertainties to the adversary, making it harder to win the distinguishing game, leading to a weaker/smaller lower bound.
What differing data $\mathbf{z}$ benefits the adversary? Compare cases where $\mathbf{z}$ is some real-world data (e.g., $\mathbf{z}=\mathbb{M}[0]$) with cases where $\mathbf{z}=\mathbf{Z}_{D}$, again, the adversary favors the contrived setting. This is because, by restricting the observation to only one dimension, the adversary avoids the uncertainties that are harder to capture under the high-dimension setting, gaining the adversary more distinguishing power. In fact, $\mathbf{z}=\mathbf{Z}_{D}$ suffices to be worst-case, as will be shown in Section 6.1 due to the rotational-invariant property of Gaussian noise.
Proxy $\varepsilon_{B}$ $\varepsilon_{U}$ $\varepsilon_{L}$
NTNV @@@@ $X=$ NTCV @@@@ $X=$ ETCV @@@@ $X=$
[…]
$\lambda_{a}$ $\lambda_{b}$ 1 1.86 $0.54$ $0.54$ $0.84$ $0.77$ $0.88$ $0.81$ $0.52$ $0.52$ $0.99$ $1.03$ $1.06$ $1.11$ $1.17$ $1.19$
$\lambda_{a}$ $\lambda_{b}$ 2 3.59 $0.99$ $1.00$ $1.41$ $1.37$ $1.35$ $1.32$ $0.91$ $0.91$ $1.17$ $1.19$ $1.17$ $1.19$ $2.36$ $2.09$
$\lambda_{a}$ $\lambda_{b}$ 4 6.79 $1.91$ $1.91$ $2.13$ $2.13$ $2.13$ $2.13$ $1.74$ $1.74$ $1.99$ $1.98$ $1.99$ $1.98$ $4.28$ $4.68$
TABLE III: Results summary for the lower bound $\varepsilon_{L}$ (averaged) obtained due to two different proxies under $\mathbf{z}=\mathbf{z}_{D}$ setup.
How does the score function (selection) leak privacy? This question is of much more interest than the previous two because it has never been studied in all previous auditing tasks (auditing the DP-SGD protocol) [39, 23, 37]. We summarise some representative results of our audit experiments in Table III. First, we observe that $\lambda_{a}$ and $\lambda_{b}$ have similar performance across all setups. Second, by comparing NTNV and NTCV, we see that manipulating the score function does not bring noticeable distinguishing advantages to the adversary.
This phenomenon may seem counter-intuitive and contrasts to the “magic” the adversary may expect: the adversary expects the (distribution of) score of the output when $\mathbf{z}$ is included in the training to be much different from that when $\mathbf{z}$ is not included; hence, manipulating the metric to select the best candidate, should make it much easier for the adversary to succeed the distinguishing game. However, our experimental results diverge from this expectation. The explanation lies in the following observations. $\lambda_{a}$ is computed based on the output resulting from selection by the score function, i.e., the selection action has already included the score information, indicating $\lambda_{b}$ is a duplicate of $\lambda_{a}$.
In principle, the score function’s output is differentially private as it is just post-processing of the trained model. With or without $\mathbf{z}$, the induced score distribution is close to each other, just as DP guarantees. Additional experimental results are presented in Appendix A.4 to validate such a claim. The “magic” cannot happen, and it holds for any score function, given that it is independent of the sensitive data.
5 Improved Privacy Results
In our previous empirical study, a conspicuous gap still exists between $\varepsilon_{U}$ derived by [42] and $\varepsilon_{L}$ obtained even under a strong adversary. Our study in the remaining sections shows such a gap exists for two reasons.
  1. 1.
    Current generic privacy upper bound is not tight;
  2. 2.
    Adversary’s power is probably not strong enough because it is hard for the adversary to instantiate the worst-case score function $g$.
Regarding 1), we provide improved privacy results and elucidate on the special property of DP-SGD leading to the improvement; for 2), we present meaningful findings about the score function.
Problem modelling. Informally, the privacy problem for our private tuning algorithm $\mathcal{H}$ (Algorithm 2) can be compactly described as the following optimization formulation.
minimize: $\varepsilon_{\mathcal{H}}$ (11)
[…]
 base algorithm’s privacy is ✦
It is self-evident that the tightness of $\varepsilon_{\mathcal{H}}$ depends on how tight ✦ is. The critical part is how we represent the privacy guarantee ✦. Under our optimization formulation, previous work describes ✦ as follows: 1) Liu et.al [30] represents the base algorithm by $(\varepsilon,\delta)$-DP; 2) Papernot et al. [42] does that by $(\alpha,\gamma)$-RDP, obtaining improved results over [30]. Can we do better? As will be shown below, the answer is yes if we represent the base algorithm’s privacy by $f$-DP.
5.1 Preliminaries: $f$-DP
$f$-DP [17], a privacy formulation with a finer resolution, reflects the nature of private mechanisms by a function [58] rather than a single pair of parameters. Our improved results are derived based on the $f$-DP framework. We introduce the necessary definitions and technical preliminaries in the following.
Definition 3 (Trade-off function [17]).
For a hypothesis testing problem over two distributions $P,P^{\prime}$, define the trade-off function as:
$T_{P,P^{\prime}}(\mathrm{FP})=\inf_{\mathcal{R}}\{\mathrm{FN}_{\mathcal{R}}:\mathrm{FP}_{\mathcal{R}}\leq\mathrm{FP}\}$
where decision rule $\mathcal{R}$ takes input a sample from $P$ or $P^{\prime}$ and decides which distribution produced that sample. The infimum is taken over all decision rule $\mathcal{R}$.
The trade-off function governs the best one can achieve when distinguishing $P$ from $P^{\prime}$, i.e., by the optimal/smallest type II error ($\mathrm{FN}$) at fixed type I error ($\mathrm{FP}$). The optimal $\mathrm{FN}$ is achieved via the likelihood ratio test, which is also known as the fundamental Neyman–Pearson lemma [41] (please refer to Appendix A.1). We denote
[…]
Definition 4 ($f$-DP [17]).
Let $f:[0,1]\rightarrow[0,1]$ be a trade-off function. A mechanism $\mathcal{M}$ satisfies f-DP if
$T_{\mathcal{M}(X),\mathcal{M}(X^{\prime})}\geq f$
for all adjacent dataset pairs $X,X^{\prime}$
$\mathcal{M}$ being $f$-DP means that any possible error pair $(\mathrm{FP},\mathrm{FN})$ resulting from distinguishing $\mathcal{M}(X)$ from $\mathcal{M}(X^{\prime})$ is lower-bounded by the curve specified by $f$. To see why $(\varepsilon,\delta)$-DP is loose. We must express $(\varepsilon,\delta)$-DP with the language of $f$-DP. This is done via the following proposition.
Proposition 1 ($(\varepsilon,\delta)$-DP equals to $f_{\varepsilon,\delta}$-DP [17, 53]).
$\mathcal{M}$ is $(\varepsilon,\delta)$-DP if and only if it is $f_{\varepsilon,\delta}$-DP where the trade-off function $f_{\varepsilon,\delta}$ is
$f_{\varepsilon,\delta}(x)=\max{(0,1-\delta-\mathrm{e}^{\varepsilon}x,\mathrm{e}^{-\varepsilon}(1-\delta-x))}$
$f$-DP implies $(\varepsilon,\delta)$-DP and conversion from $f$-DP to $(\varepsilon,\delta)$-DP is via Algorithm 3 (restatement of Proposition 6 of [17]).
Algorithm 3 $f$-DP to $(\varepsilon,\delta)$-DP
1:trade-off function $f$, $\delta$ with $\delta\geq 1-f(0)$
2:Compute $\varepsilon=\inf\{a:f(x)\geq 1-\delta-\mathrm{e}^{a}x,\forall x\in[0,1]\}$ via binary search
3:$\max\{0,\varepsilon\}$
In plain words, $f_{\varepsilon,\delta}$-DP (or $(\varepsilon,\delta)$-DP) for some mechanism $\mathcal{M}$ is the two symmetric straight lines lower-bounding the true/faithful trade-off function of $\mathcal{M}$. This is drawn in Figure 6. For the Gaussian mechanism, which is probably the most basic private mechanism, using $(\varepsilon,\delta)$-DP to characterize its privacy is not tight/faithful; in contrast, the following special family of trade-off functions is tight.
Definition 5 ($\mu$-Gaussian DP ($\mu$-GDP) [17]).
The trade-off function of distinguishing $\mathcal{N}(0,1)$ from $\mathcal{N}(\mu,1)$ is
$G_{\mu}(x)=T_{\mathcal{N}(0,1),\mathcal{N}(\mu,1)}(x)=\Phi(\Phi^{-1}(1-x)-\mu),$
where $\Phi$ be the c.d.f. of standard normal distribution. A private mechanism $\mathcal{M}$ satisfies $\mu$-GDP if it is $G_{\mu}$-DP
The analytical expression of $\mu$-GDP is determined by applying the Neyman–Pearson lemma on the hypothesis testing problem of distinguishing $\mathcal{N}(0,1)$ from $\mathcal{N}(\mu,1)$ [17]. $\mathcal{M}$ satisfying $\mu$-GDP means that distinguishing $\mathcal{M}(X)$ from $\mathcal{M}(X^{\prime})$ is at least as hard as distinguishing two univariate Gaussians $\mathcal{N}(0,1),\mathcal{N}(\mu,1)$ based on a single draw. Figure 6 explains why $(\varepsilon,\delta)$-DP is loose: $(\varepsilon,\delta)$-DP is strictly more conservative than $\mu$-GDP when characterizing the privacy of Gaussian mechanism.
Figure 6: For the Gaussian mechanism $\mathcal{M}(X)=q(X)+\mathcal{N}(0,\sigma^{2}\mathbb{I}^{d})$ where the query function $q(X)\in\mathbb{R}^{d}$ has unit $\ell_{2}$-sensitivity, it is exactly $1/\sigma$-GDP [17]. It is also some $(\varepsilon,\delta)$-DP; however, $\mu$-GDP characterization saves the shaded area in gray.
The following corollary gives a closed-form solution for optimal/lossless conversion from $\mu$-GDP to $f_{\varepsilon,\delta}$-DP (or $(\varepsilon,\delta)$-DP) in accordance with Algorithm 3.
Corollary 1 (Conversion from $\mu$-GDP to $(\varepsilon,\delta)$-DP formulation [17, 5]).
A mechanism is $\mu$-GDP if and only if it is $f_{\varepsilon,\delta(\varepsilon)}$-DP (or $(\varepsilon,\delta(\varepsilon))$-DP) $\forall\varepsilon\geq 0$ where
$\delta(\varepsilon)=\Phi(-\frac{\varepsilon}{\mu}+\frac{\mu}{2})-\mathrm{e}^{\varepsilon}\Phi(-\frac{\varepsilon}{\mu}-\frac{\mu}{2})$ (12)
Remark 1.
The purpose of introducing all previous technical preliminaries (especially Figure 6) is not only to necessarily introduce $f$-DP itself but also to understand why we can obtain improvements under $f$-DP framework (see Example 2).
DP-SGD is asymptotically $\mu$-GDP. The pivotal role of $\mu$-GDP is that it asymptotically characterizes the privacy of any DP-SGD instance having many compositions of Gaussian mechanisms [17], as shown in the following.
Corollary 2 (GDP approximation [17] for DP-SGD).
DP-SGD is asymptotically $\mu$-GDP with
$\mu=\sqrt{2}\tau\sqrt{N}\cdot\sqrt{\mathrm{e}^{\sigma^{-2}}\cdot\Phi\left(1.5\sigma^{-1}\right)+3\Phi\left(-0.5\sigma^{-1}\right)-2}$
where $\sigma=\sigma^{\prime}/C$ and $\sigma^{\prime}$ is s.t.d. of the Gaussian noise; $C$ is the clipping threshold; $\tau$ and $N$ is the sampling ratio and number of total iteration of DP-SGD.
5.2 Our Contribution: Improved Results
For Corollary 2, the error (pointwise error between the asymptotical GDP trade-off function and the true trade-off function) decays at a rate of $1/\sqrt{N}$ for DP-SGD, as shown by the analysis in [17]. Therefore, using Corollary 2 requires $N$ to be large enough. Such a condition holds for probably most DP-SGD applications, especially for training large models (e.g., $N>10^{4}$ in [2] and $N>10^{5}$ in [27, 51]). Based on all of the above preparations, we are ready to approach our privacy problem by filling in the missing part of Equation (11) based on Corollary 2:
minimize: $\varepsilon_{\mathcal{H}}$ (13)
[…]
 base algorithm’s privacy is $\mu$-GDP
In the following, we first revisit the central question of how selection (the score function) leaks privacy.
Recall in Section 4.6, we showed that the adversary gains no advantage even if the score function $g$ is maliciously manipulated. A natural question arises: is there any score function that brings more advantage to the adversary? The answer is positive as there is a special type of score function that tends to leak more sensitive information, regardless of whether manipulated or not.
One-to-one mapping $g$ is the worst-case necessarily. Before introducing our main privacy result, we have a final question remaining to be answered: how does the score function $g$ affect the privacy of $\mathcal{H}$? Specifically, as a function that maps the output of the base algorithm to a real number, If some $g$ happens to map two distinct inputs to the same score (hence, a randomized tie-breaking will be enforced), how does such $g$ affect the privacy of $\mathcal{H}$ compared to one-to-one mapping score functions?
Intuitively, such $g$ will only make $\mathcal{H}$ more private as new uncertainty is injected. We can gain more intuition by considering the extreme case: if $g$ only outputs a constant, then $\mathcal{H}$ is just as private as the base algorithm. Our theorem in the following formalizes such intuition.
Theorem 2 (Necessary worst-case $g$, proof in Appendix B.2).
Let distribution $P$ be over some finite alphabets $\Gamma$, and define a distribution $F_{k,g}$ as follows. First, make $k>0$ independent samples $\{x_{1},x_{2},\cdots,x_{k}\}$ from $P$; second, output $x_{i}$ such that the score $g(x_{i})$ computed by a score function $g:\Gamma\rightarrow\mathbb{R}$ is the maximum over these samples. Similarly, we define another distribution $P^{\prime}$ over the same alphabets $\Gamma$ and derive a distribution $F^{\prime}_{k,q}$ as the counterpart to $F_{k,g}$.
For any score function $\hat{g}$, which is not a one-to-one mapping (hence a randomized tie-breaking is needed), there always exists a one-to-one mapping $g^{*}$ satisfying
$\mathcal{D}_{\alpha}(F_{k,\hat{g}}||F^{\prime}_{k,\hat{g}})\leq\mathcal{D}_{\alpha}(F_{k,g^{*}}||F^{\prime}_{k,g^{*}}).$ (14)
Moreover, similar inequality also holds when $k$ follows a general distribution $\xi$.
The above result is derived under RDP (Definition 2) due to analytical feasibility, but it tells us crucial facts: A score function that induces a strict total order for elements in $\Gamma$ tends to be less private. Thus, a one-to-one mapping is necessary to be the worst case for the score function $g$. Theorem 2 also holds when $\Gamma$ is infinite because Rényi divergence can be approximated arbitrarily well by finite partition [50, Theorem 10].
With $g$’s necessary condition determined, we formalize some notation and introduce our improved privacy results in the following.
Notation. Let $y,y^{\prime}\in\mathcal{Y}$ be the output of the base algorithm (DP-SGD, a single run) corresponding to adjacent input dataset $X,X^{\prime}$, respectively. Let $P,P^{\prime}$ be the induced score distribution after the score function $g$ takes input $y,y^{\prime}$, respectively. With some abuse of notation, we use $P(x),F(x)$ to denote the p.d.f. and c.d.f. for distribution $P$ (similarly, we have $P^{\prime}(x),F^{\prime}(x)$ w.r.t. $X^{\prime}$). Based on the assumption that $g$ is a one-to-one mapping, the selection is essentially among samples from $P$ (or $P^{\prime}$ if $X^{\prime}$ is the input).
Let $Q$ be the distribution of the score of the model outputted by $\mathcal{H}$. Let us for now consider the distribution $\xi$ in $\mathcal{H}$ is a point mass on some $k>0$, i.e., $\operatorname{Pr}(k)=1$. Then, the p.d.f. $Q(x)$ is
$Q(x)=kP(x)(F(x))^{k-1}$ (15)
as well-studied in order statistics [14], i.e., it is the distribution of the maximal sample among $k$ independent draws.
When distribution $\xi$ is some general distribution, define the function
$\omega_{\xi}(x)=\sum_{k\sim\xi}k\cdot\operatorname{Pr}_{\xi}(k)\cdot x^{k-1}$ (16)
and then $Q$ is a mixture distribution, i.e.,
$Q(x)=\sum_{k\sim\xi}\operatorname{Pr}_{\xi}(k)\cdot kP(x)(F(x))^{k-1}=P(x)\omega_{\xi}(F(x)).$ (17)
Distribution $Q^{\prime}$’s p.d.f. corresponding to $X^{\prime}$ being the input is computed similarly. Now, we are ready to present our improved privacy upper bound.
Theorem 3 (General form, proof in Appendix B.3).
Suppose the base algorithm is $f$-DP, then $\mathcal{H}$ is $(\varepsilon_{\mathcal{H}},\delta_{\mathcal{H}})$-DP where
$\varepsilon_{\mathcal{H}}=\varepsilon+\max_{a\in[0,1]}\log{{\frac{\omega_{\xi}(1-a)}{\omega_{\xi}(b)}}},$ (18)
where $b=f(a)$ and $\varepsilon$ is computed by Algorithm 3 whose two input arguments are the trade-off function $f$ and $\delta=\delta_{\mathcal{H}}/\omega_{\xi}(1)$ ($\omega_{\xi}$ is defined in Equation (16)).
We present our $f$-DP accountant for private selection in Algorithm 4 according to Theorem 3.
Algorithm 4 $f$-DP Accountant for $\mathcal{H}$
1:trade-off function $f$ s.t. the base algorithm is $f$-DP, $\xi$ distribution of $\mathcal{H}$, $\delta_{\mathcal{H}}$
2:$\delta\leftarrow\delta_{\mathcal{H}}/\omega_{\xi}(1)$$\triangleright$ $\omega_{\xi}$ is from Equation (16)
3:$\varepsilon\leftarrow$ input $f$ and $\delta$ to Algorithm 3
4:$\varepsilon_{\mathcal{H}}\leftarrow\varepsilon+\max_{a\in[0,1]}\log{{\frac{\omega_{\xi}(1-a)}{\omega_{\xi}(f(a))}}}$
5:$\varepsilon_{\mathcal{H}}$
Given that the base algorithm is some $\mu$-GDP, we immediately arrive at the improved result for hyper-parameter tuning by plugging in its specific trade-off function.
Corollary 3 (Improved result for DP-SGD).
If the base algorithm if $\mu$-GDP (or $G_{\mu}$-DP), then $\mathcal{H}$ is $(\varepsilon_{\mathcal{H}},\delta_{\mathcal{H}})$-DP where
$\varepsilon_{\mathcal{H}}=\varepsilon+\max_{a\in[0,1]}\log{{\frac{\omega_{\xi}(1-a)}{\omega_{\xi}(G_{\mu}(a))}}}$ (19)
with $G_{\mu}(a)$ is in Definition 5 and $\delta_{\mathcal{H}}/\omega_{\xi}(1)=\Phi(-\frac{\varepsilon}{\mu}+\frac{\mu}{2})-\mathrm{e}^{\varepsilon}\Phi(-\frac{\varepsilon}{\mu}-\frac{\mu}{2})$ determines $\varepsilon$.
Why we can obtain improvements. In the following, we show why modeling the base algorithm with $f$-DP other than $(\varepsilon,\delta)$-DP brings improvement. By post-processing property [17], the score outputted by the score function satisfies the same $f$-DP as that of the base algorithm. Term $a$ in Equation (18) can be seen as the $\mathrm{FN}$ for distinguishing $P$ from $P^{\prime}$ and $b=f(a)$ is the optimal $\mathrm{FP}$ at $\mathrm{FN}=a$.
For DP-SGD instance satisfying $\mu$-GDP, we have $b=G_{\mu}(a)$, however, if its privacy is represented by some $(\varepsilon,\delta)$-DP (or $f_{\varepsilon,\delta}$-DP) we can only get $b=f_{\varepsilon,\delta}(a)$ in Equation (18). Note that figure 6 tells us that $G_{\mu}(a)\geq f_{\varepsilon,\delta}(a)$, hence modeling the base algorithm by $(\varepsilon,\delta)$-DP only leads to higher (weaker) $\varepsilon_{\mathcal{H}}$ value in Equation (18) as $\omega_{\xi}$ is increasing. The following numerical example demonstrates the superiority of modeling within the $f$-DP framework.
Example 2.
Suppose the base algorithm (DP-SGD) satisfies $1$-GDP and $\xi$ is the TNB distribution with parameter $\eta=1,\nu=10^{-2}$ (in this case, $\xi$ is geometric distribution, and we recover the case studied by Liu et al. [30]). Hence, it allows us to make meaningful comparisons.
For $\delta=10^{-5}$, the base algorithm is also $(4.36,10^{-5})$-DP or $f_{4.36,10^{-5}}$-DP. If $b=G_{1}(a)$ in Equation (18), which is how we represent the base algorithm’s privacy, we have $\max_{a\in[0,1]}\log{{\frac{\omega_{\xi}(1-a)}{\omega_{\xi}(G_{1}(a))}}}=3.3$; however, if $b=f_{4.36,10^{-5}}(a)$, which equals to how the base algorithm is modeled by Liu et al. [30], we can only have $\max_{a\in[0,1]}\log{{\frac{\omega_{\xi}(1-a)}{\omega_{\xi}(f_{4.36,10^{-5}}(a))}}}=16.5>3.3$. Thus, a huge improvement is obtained, and this is due to the saved shaded area in gray shown in Figure 6.
5.3 Discussion
Improvement is also possible in general. Our above example shows that representing the privacy of the base algorithm with finer resolution (from $(\varepsilon,\delta)$-DP to $f$-DP) leads to improvements in the privacy result for the private selection problem. Similar conclusions also hold when switching from RDP [42] to $f$-DP as RDP is also observed to be lossy within the $f$-DP framework [58], i.e., RDP shares the same weakness as that of the $(\varepsilon,\delta)$-DP. Therefore, for any other private mechanisms analyzed by $(\varepsilon,\delta)$-DP or RDP, switching to $f$-DP also brings improvements possibly.
Our results are both tighter and more generalizable. As will be shown in Section 6.2, our result is tighter. Moreover, consider a $\xi$ distribution such that $\operatorname{Pr}_{\xi}(1)=1$ (i.e., run the base algorithm only once), we see that Equation (18) becomes $\varepsilon_{\mathcal{H}}=\varepsilon$ and $\delta_{\mathcal{H}}=\delta$ for any $(\varepsilon,\delta)$ pair, which is the base algorithm’s privacy guarantee. This shows that our result is tight for general distribution $\xi$ in terms of how much $(\varepsilon_{\mathcal{H}},\delta_{\mathcal{H}})$-DP $\mathcal{H}$ satisfies. In contrast, analysis by RDP can’t achieve such tightness as converting RDP to $(\varepsilon,\delta)$-DP is lossy [58].
As for the generalizability, we note that 1) [30] only works with $\xi$ being the geometric distribution; 2) [42] only works with $\xi$ being the TNB family and Poisson distribution. It is unknown how to handle the case where $\xi$ is some arbitrary distribution, and it seems like at least case-by-case analysis is needed if using previous approaches. In contrast, our result applies to any distribution $\xi$ of protocol $\mathcal{H}$. Also, note that computing $\omega_{\xi}$ for any reasonable distribution is always numerically stable as $\omega_{\xi}(x)$ is bounded in $[0,1]$.
6 Further Evaluation
Unlike the experiments in Section 4, which relies on some ad hoc audit setups, our further audit experiments in this section make reductions to the audit and confine to the idealized settings of the threat model of DP. We aim to explore the tightness of our improved result by finding where the privacy lower bound is under an adversary with compelling and theoretically justified power.
6.1 Stronger Audit via Reduction
1) Base algorithm reduction. To have stronger privacy audit result over $\mathcal{H}$, it is imperative to first have stronger audit on $\mathcal{H}$’s the base DP-SGD algorithm. This part is devoted to this topic.
Our threat model will be based on the assumption made by DP, i.e., the adversary knows all but one data, a.k.a. the strong adversary assumption [13, 29]. Translated to our setting, the adversary knows the membership of all data used to update the model in each iteration except for the differring data $\mathbf{z}$. This means the adversary can always subtract the gradient of other data from $p_{i}$ in each iteration. Based on such setups, any adjacent dataset pair $X,X^{\prime}$ is equivalent to $X=\emptyset,X^{\prime}=\mathbf{z}$ from the adversary’s view, i.e., we audit in an idealized setting.
The remaining part is how we form a compelling proxy and let the adversary make assertions based on it. This is done via two reductions for the base algorithm (DP-SGD).
❶ First reduction (from Equation (20) to Equation (22)). Let $\sigma$ noise s.t.d. shown in Equation (1). Now assume that the sampling ratio $\tau=1$, i.e., full-batch gradient descent. Given $X=\emptyset,X^{\prime}=\mathbf{z}$, then, at each iteration, for the adversary, the private gradient $p_{i}$ is as follows.
$p_{i}|X=R_{i}\sim\mathcal{N}(0,C^{2}\sigma^{2}\mathbb{I}^{d})$ (20)
$p_{i}|{X^{\prime}}=(\nabla_{\mathbf{z}}+R_{i})\sim\mathcal{N}(\nabla_{\mathbf{z}},C^{2}\sigma^{2}\mathbb{I}^{d}),$
where $p_{i}|X$ denotes the random variable conditioned on $X$ was chosen and $\nabla_{\mathbf{z}}=\nabla_{w}\ell(w_{i-1},z)$ with $\|\nabla_{\mathbf{z}}\|_{2}=C$ (assume maximal $\ell_{2}$-norm is reached). The adversary can always construct a rotational matrix $U_{\mathbf{z}}\in\mathbb{R}^{d\times d}$ such that $p_{i}$ can be reduced as follows.
$U_{\mathbf{z}}p_{i}|X\sim\mathcal{N}(0,C^{2}\sigma^{2}\mathbb{I}^{d}),\text{\quad}U_{\mathbf{z}}p_{i}|X^{\prime}\sim\mathcal{N}(U_{\mathbf{z}}\nabla_{\mathbf{z}},C^{2}\sigma^{2}\mathbb{I}^{d})$ (21)
where $U_{\mathbf{z}}\nabla_{\mathbf{z}}=[C,0,0,\cdots]^{T}$. This is because Gaussian noise with $\sigma^{2}\mathbb{I}^{d}$ covariance is rotational invariant, i.e.,
$\mathbf{Cov}(U_{\mathbf{z}}R_{i})=U_{\mathbf{z}}\mathbf{Cov}(R_{i})U_{\mathbf{z}}^{T}=C^{2}\sigma^{2}\mathbb{I}^{d}=\mathbf{Cov}(R_{i})$
where $\mathbf{Cov}(R_{i})$ is the covariance matrix of Gaussian random vector $R_{i}$. After the rotation, for the adversary, only the first coordinate carries useful information about $\mathbf{z}$.
This is because a noise vector $R_{i}\sim\mathcal{N}(0,C^{2}\sigma^{2}\mathbb{I}^{d})$ and its rotated version $U_{\mathbf{z}}R_{i}\sim\mathcal{N}(0,C^{2}\sigma^{2}\mathbb{I}^{d})$ possessing $\sigma^{2}\mathbb{I}^{d}$ covariance matrix are all coordinate-wise independent. To serve the distinguishing purpose: $\mathbf{z}$ was/was not used, it suffices to characterize the private gradient $p_{i}$ by $\bar{p}_{i}$ as a univariate random variable (the first coordinate) for the distinguishing purpose as follows.
$\bar{p}_{i}|X\sim\mathcal{N}(0,C^{2}\sigma^{2}),\text{\quad}\bar{p}_{i}|X^{\prime}\sim\mathcal{N}(C,C^{2}\sigma^{2}).$ (22)
Therefore, our above analysis justifies the sufficiency of the Dirac canary $\mathbf{z}=\mathbf{Z}_{D}$ being worst-case in our previous experiments in Section 4.5.
❷ Second reduction (from Equation (22) to Equation (24)). By the first reduction and from the adversary’s view, the DP-SGD’s output $\bar{y}=\{\bar{p}_{1},\bar{p}_{2},\cdots,\bar{p}_{N}\}$ is essentially an observation of $N$ i.i.d. samples from $\mathcal{N}(a,C^{2}\sigma^{2})$ where $a$ is either 00 or $C$. Recall the adversary’s goal is to distinguish $X$ or $X^{\prime}$ was used; this is equivalent to determining $a=0$ or $a=C$.
As both distributions in Equation (22) are Gaussian, we can use the sufficient statistics for estimating $a$ [21, 12], which is the mean of such $N$ i.i.d. samples: $\bar{y}=\frac{1}{N}\sum_{i=1}^{N}\bar{p}_{i}$. Sufficient statistics do not lose any information for estimating $a$.
Finally, we can reduce the privacy of the base algorithm to an equivalent game for the adversary as
$\bar{y}|X\sim\mathcal{N}(0,C^{2}\sigma^{2}/N),\text{\quad}\bar{y}|X^{\prime}\sim\mathcal{N}(C,C^{2}\sigma^{2}/N).$ (23)
For simplicity, applying a simple invertible/lossless re-scaling gives us equivalent characterization:
$\bar{y}|X\sim\mathcal{N}(0,1),\text{\quad}\bar{y}|X^{\prime}\sim\mathcal{N}(\sqrt{N}/\sigma,1).$ (24)
There is a slight difference in the reduction when $\tau<1$. Instead of arriving at Equation (22), we arrive at
$\bar{p}_{i}|X\sim\mathcal{N}(0,C^{2}\sigma^{2}),\text{\quad}\bar{p}_{i}|X^{\prime}\sim\mathcal{N}(Cb_{i},C^{2}\sigma^{2}),$ (25)
where $b_{i},\forall i\in\{1,2,\cdots,N\}$ is independent Bernoulli random variables with probability $\tau$. By doing the same transformation as from Equation (22) to Equation (24), we arrive at
$\bar{y}|X\sim\mathcal{N}(0,1),\text{\quad}\bar{y}|X^{\prime}\sim\mathcal{N}(\frac{\sum_{i=1}^{N}b_{i}}{N}\sqrt{N}/\sigma,1).$ (26)
Equation (26) also covers Equation (24) when $\tau=1$.
Figure 7: The distinguishing game for $\mathcal{H}$ with the base algorithm reduced. Two univariate Gaussians $P=\mathcal{N}(0,1)$ and $P^{\prime}=\mathcal{N}(\frac{\sum_{i=1}^{N}b_{i}}{N}\sqrt{N}/\sigma,1)$ correspond to adjacent dataset pair $X,X^{\prime}$. At each execution, the trainer chooses one distribution and feeds it to $\mathcal{H}$; then an integer $k>0$ is sampled from $\xi$ followed by independently sampling $k$ times from $\hat{P}$. After getting such $k$ samples, the best sample $\bar{y}_{i}$ is selected based on a score function $g(x)=x$. The adversary then makes binary assertions by 1) simply taking the returned best sample as the proxy and 2) comparing it to some threshold. The game $\mathcal{G}$ is simulated $10^{7}$ times.
$\varepsilon_{B}$ $\tau$ $\varepsilon_{L}$ | $\varepsilon_{\mathcal{H}}^{O}$ | $\varepsilon_{\mathcal{H}}^{P}$
$\eta=0,\nu=10^{-2}$ $\mathbb{E}_{\xi}=21$ $\eta=1,\nu=10^{-2}$ $\mathbb{E}_{\xi}=100$ $\eta=1,\nu=10^{-3}$ $\mathbb{E}_{\xi}=1000$ $\eta=2,\nu=10^{-3}$ $\mathbb{E}_{\xi}=2000$
1 1 $1.17$ | $1.55$ | $1.86$ $1.19$ | $2.06$ | $2.65$ $1.27$ | $2.54$ | $3.09$ $1.29$ | $3.18$ | $3.99$
0.5 $1.09$ | $1.53$ | $1.87$ $1.30$ | $2.04$ | $2.64$ $1.37$ | $2.51$ | $3.15$ $1.42$ | $3.15$ | $4.05$
0.1 $1.00$ | $1.49$ | $1.86$ $1.19$ | $1.98$ | $2.56$ $1.49$ | $2.43$ | $3.24$ $1.46$ | $3.05$ | $4.09$
2 1 $2.17$ | $2.92$ | $3.61$ $2.21$ | $3.84$ | $5.06$ $2.53$ | $4.69$ | $5.89$ $2.57$ | $5.85$ | $7.57$
0.5 $2.05$ | $2.89$ | $3.57$ $2.34$ | $3.80$ | $4.93$ $2.59$ | $4.64$ | $5.88$ $2.63$ | $5.79$ | $7.49$
0.1 $2.16$ | $2.90$ | $3.58$ $2.40$ | $3.82$ | $4.83$ $2.87$ | $4.67$ | $6.03$ $2.86$ | $5.82$ | $7.53$
4 1 $3.93$ | $5.70$ | $6.80$ $4.32$ | $7.40$ | $9.30$ $4.51$ | $8.95$ | $10.83$ $4.70$ | $11.07$ | $13.77$
0.5 $3.84$ | $5.64$ | $6.71$ $4.23$ | $7.32$ | $9.03$ $4.89$ | $8.86$ | $10.74$ $4.91$ | $10.96$ | $13.51$
0.1 $3.76$ | $5.48$ | $6.58$ $4.17$ | $7.12$ | $8.64$ $5.02$ | $8.62$ | $10.61$ $5.03$ | $10.67$ | $13.08$
TABLE IV: Comparison of various privacy results for $\mathcal{H}$ (Algorithm 2). All results are in $(\varepsilon,\delta=10^{-5})$-DP form. Value for $\varepsilon_{L}$ (the lower bound) is the average of 10 runs with different seeds for simulating the game shown in Figure 7. $\varepsilon^{O}_{\mathcal{H}}$ is our improved upper bound (Algorithm 4 with the base algorithm is some $\mu$-GDP) and $\varepsilon_{\mathcal{H}}^{P}$ is the privacy upper bound derived by previous work [42]. $\varepsilon_{L}$, $\varepsilon^{O}_{\mathcal{H}}$ and $\varepsilon_{\mathcal{H}}^{P}$ in each cell are all computed based on the same base algorithm’s privacy budget $\varepsilon_{B}$ shown in leftmost column. Note that $\tau$ is the sampling ratio with $N=10^{3}$ total iteration, and $\eta,\nu$ is the parameter for TNB distribution (shown in Appendix A.2; $\mathbb{E}_{\xi}$ is its expectation) input to $\mathcal{H}$. We can see that our result is better under all setups.
2) Instantiate the score function. Based on our above reduction, to serve the distinguishing purpose, a model obtained by DP-SGD can be “treated” as a real number sampled from univariate Gaussian or its shifted counterpart corresponding to $X$ or $X^{\prime}$ was used. With our reduction, the order induced by the score function $g$ is now over $\mathbb{R}$. To have stronger audit results in our idealized attack, we need to instantiate the worst-case score function, and our Theorem 2 tells us one-to-one mapping score function $g$ is the worst case necessarily.
However, Theorem 2 remains silent on the specific analytical form of $g$ in the worst case. We will discuss the hardness of finding the specific worst-case $g$ in Section 6.2. Based on our above reduction, there can be infinitely many one-to-one mapping functions $\mathbb{R}\rightarrow\mathbb{R}$; for implementation purposes, we now fix a score function $g(x)=x$, i.e., we take $g$ is strictly increasing and all strictly increasing functions induces the same order over $\mathbb{R}$ regardless of its analytical form. To summarise, we present the distinguishing game of our reduced case in Figure 7. Our further audit experiment in this section is to simulate such a game many times and conclude the lower bound in accordance with Section 2.2.
6.2 Lower Bound & Privacy Results Comparison
We present the lower bound $\varepsilon_{L}$, our improved privacy results $\varepsilon^{O}_{\mathcal{H}}$ and previous privacy results $\varepsilon_{\mathcal{H}}^{P}$ [42] in Table IV. Fixing $\delta_{\mathcal{H}}=10^{-5}$, we can see that our privacy result $\varepsilon_{\mathcal{H}}^{O}$ is better than previous result $\varepsilon_{\mathcal{H}}^{P}$ under all setups. We can also see that $\varepsilon_{L}$ is higher than $\varepsilon_{B}$, confirming the very action of selecting the best does leak additional privacy beyond the base algorithm’s privacy budget, similar to our finding in Section 4.5 (ETCV). However, we must emphasize that such a phenomenon is only observed under a strong adversary setup.
Lessons Learned and Open Problems. As shown in Section 5.3, our improved result is indeed tight for general $\xi$ in terms of how much $(\varepsilon_{\mathcal{H}},\delta_{\mathcal{H}})$-DP $\mathcal{H}$ satisfies. However, we still see a noticeable gap between our result $\varepsilon^{O}_{\mathcal{H}}$ and the lower bound $\varepsilon_{L}$ derived by our idealized attack. why does this happen? Answering this question will lead us to a deeper understanding of how selection leaks privacy.
Our answer is that $g$ plays a critical role from the adversary’s point of view, and such a factor distinguishes attacking private hyper-parameter tuning from all previous privacy attack problems. We have shown that one-to-one mapping is necessary for $g$ being the worst case, but there are infinitely many $g$ over an infinite output domain and are up to $|\Gamma|!$ possible choices if output domain $\Gamma$ is finite. We find that some of the score functions leak more privacy than others. For example, when $\eta=1,\nu=10^{-2}$ (TNB $\xi$ recovers geometric distribution), if we (arbitrarily) set the score function as
$g(x)=\left\{\begin{array}[]{ll}x,\hskip 15.93347pt\text{for $x\in[-\infty,0)\bigcup(1,\infty]$}\\ 1-x,\text{for $x\in[0,1]$}\\ \end{array}\right.$ (27)
which is clearly a one-to-one mapping, we only derive $\varepsilon_{L}=2.01$ (average of 10 runs) at $\varepsilon_{B}=2,\tau=1$, which is smaller/weaker than the value $2.21$ shown in Table IV (where $g(x)=x$).
Choosing some $g$ arbitrarily and performing the attack will likely end up with sub-optimal attacks (smaller/weaker lower bounds). This is possibly the reason why we still see a gap between $\varepsilon^{O}_{\mathcal{H}}$ and $\varepsilon_{L}$ even in our idealized attack. This poses a critical question for the adversary, i.e., which $g$ should the adversary choose to elicit more privacy leakage? Another equally important question is: does the worst-case $g$ depend on specific $\xi$? Answering the above questions is non-trivial and requires future investigations.
7 Conclusion
We aim to gain a deeper understanding of current theoretical v.s. practical privacy for the private selection problems. Initially, we give examples showing that the current generic bound for private selection is indeed tight in general. However, when we consider a white-box setting, i.e., the hyper-parameter tuning problem, such tightness no longer applies. Substantiating this assertion, we first conduct empirical studies for private hyper-parameter tuning under various privacy audit setups. Our findings reveal that, in certain practical scenarios, a weak adversary can only acquire restricted sensitive information. The derived empirical privacy bound for the strongest adversary still sees a noticeable gap from the generic upper bound.
We then take further steps to analyze privacy in hyper-parameter tuning and derive better privacy results by modeling the base algorithm’s privacy with finer resolution ($f$-DP). The improvement is due to the distinct properties of the base algorithm (DP-SGD). Our result is better under various settings than the current generic bound. In addition, our analysis generalizes, contrasting with previous work, where their results are only easily applicable to specific parameter configurations.
Appendix A Content for reference
A.1 Neyman–Pearson Lemma
Theorem 4 (Neyman–Pearson lemma [41]).
Let $P$ and $Q$ be probability distributions on $\Omega$ with densities $p$ and $q$, respectively. Define $L(x)=\frac{p(x)}{q(x)}$. For hypothesis testing problem
$\mathbf{H_{0}}:P,\text{\quad}\mathbf{H_{1}}:Q$
For a constant $c>0$, suppose that the likelihood ratio test which rejects $\mathbf{H_{0}}$ when $L(x)\leq c$ has $\mathrm{FP}=a$ and $\mathrm{FN}=b$, then for any other test of $\mathbf{H_{0}}$ with $\mathrm{FP}\leq a$, the achievable $\mathrm{FN}$ is at least $b$.
Neyman–Pearson lemma says that the most powerful test (optimal $\mathrm{FN}$) at fixed $\mathrm{FP}$ is the likelihood ratio test. Applying Neyman–Pearson lemma to distinguishing $\mathcal{N}(0,1)$ from $\mathcal{N}(\mu,1)$ gives us Definition 5 [17].
A.2 Privacy Results by Papernot et al.
Truncated negative binomial (TNB) distribution. For $\nu\in(0,1)$ and $\eta\in(-1,\infty)$, the distribution $\xi_{\eta,\nu}$ on $\{1,2,3,\cdots\}$ is as follows. When $\eta\neq 0$, then
$\forall k\in\mathbb{N},\quad\operatorname{Pr}[K=k]=\frac{(1-\nu)^{k}}{\nu^{-\eta}-1}\cdot\prod_{\ell=0}^{k-1}\left(\frac{\ell+\eta}{\ell+1}\right),$
when $\eta=0$, then
$\forall k\in\mathbb{N},\quad\operatorname{Pr}[K=k]=\frac{(1-\nu)^{k}}{k\cdot\log(1/\nu)}$
This particular distribution is obtained by differentiating the probability generating function of some desired form [42]. The main relevant privacy results in [42] are provided in the following. Note that they are all in RDP form.
Theorem 5 (RDP for TNB distribution [42]).
Let $k$ in Algorithm 1 follows TNB distribution $\xi_{\eta,\nu}$. If the base algorithm satisfies $(\alpha,\gamma)$-RDP and $(\alpha^{\prime},\gamma^{\prime})$-RDP, Algorithm 1 satisfies $(\alpha,\hat{\gamma})$-RDP where
$\hat{\gamma}=\gamma+(1+\eta)\cdot\left(1-\frac{1}{\alpha^{\prime}}\right)\gamma^{\prime}+\frac{(1+\eta)\cdot\log(1/\nu)}{\alpha^{\prime}}+\frac{\log\mathbb{E}_{\xi_{\eta,\nu}}}{\alpha-1}$
Tight example for approximate DP. A Tight example for $(\varepsilon,\delta)$-DP case can be obtained trivially based on Example 1. If an algorithm is $(\varepsilon,0)$-DP, it is also $(\varepsilon,\delta)$-DP. Hence, the above tight example covers the $(\varepsilon,\delta)$-DP case. Specifically, it can be checked that the example shown in Equation (5) is $(1,10^{-5})$-DP and Equation (7) is $(2.92,10^{-5})$-DP. Compared to the result predicted by [42], which is $(3.11,10^{-5})$-DP, i.e., it is tight up to a negligible gap.
A.3 Used Datasets and Experimental Details
Our implementation is provided at an anonymous link111 https://anonymous.4open.science/. We use four image datasets in our experiments. FASHION [55], MNIST [28], CIFAR10 [26] and SVHN [40]. All of our experiments are conducted under privacy parameter $\delta=10^{-5}$. The number of repeating/simulation times in an audit experiment is $2,000$. The error bar results from taking the min., max., and avg. for three trials. To efficiently audit the hyper-parameter tuning and reduce the simulation burden, we only fetch 5,000 data examples from the original training datasets, and we set the sampling rate to be 1, i.e., full gradient descent. We set the TNB distribution [42] with parameter $(\eta=0,\nu=10^{-2})$. We use the ResNet [22] as the neural network in our experiments. Details for network setup can be found in our implementations at our anonymous link. Our audit experiment consumes $>4000$ GPU hours and is conducted over 20 GPUs in parallel.
Hyperparameter candidates setup. To run the audit experiments, we need to set the candidates inside $\Omega$ in Algorithm 2. We hold the clipping threshold $C$, learning rate $\mathbf{lr}$, and the number of total iterations $N$ as the hyperparameters to be tuned. To form each candidate inside $\Omega$, we randomly sample a value to determine $C$, $\mathbf{lr}$ and $N$. Details can be found at our implementation. All candidates have the same privacy budget according to our problem formulation.
Detailed procedure of concluding the lower bound $\varepsilon_{L}$. The following procedure is adopted to conclude a lower bound $\varepsilon_{L}$.
1) Generating $(b_{\text{truth}},b_{\text{guess}})$. Each pair corresponds to an execution of $\mathcal{G}$ (Algorithm 1). The adversary needs to make an assertion, i.e., output a $b_{\text{guess}}\in\{0,1\}$.
2) Compute $\varepsilon_{L}$. After getting many pairs of $(b_{\text{truth}},b_{\text{guess}})$, the $\mathrm{FP},\mathrm{FN}$ can be summarised by Clopper-Pearson with a confidence $c$. Specifically, the $\mathrm{FP}$ rate and $\mathrm{FN}$ rate are modeled as the unknown success probabilities of two binomial distributions. Then $\varepsilon_{L}$ can be computed by Equation (4) or by the methods used in [37].
3) Optimization. In practice, practitioners often try various assertion strategies on the same observed output by repeating procedures 1) and 2) to find the optimal $\varepsilon_{L}$.
Figure 8: The output’s score histogram under all setups.
A.4 More Experimental Results
Measuring the score distribution. To confirm the robustness mentioned above. A straightforward way is to inspect the scores of the output of $\mathcal{H}$ (Algorithm 2). We measure various training datasets, and results are presented in Figure 8. We process the scores by shifting so all scores are $<0$. Our result clearly shows no significant difference between the score distributions when $\mathbf{z}$ is included ($X^{\prime}$ is used) and when it is not included ($X$ is used) in training. In Figure 8, we again obverse the “indistinguishability” between cases where $\mathbf{z}$ is included and cases where $\mathbf{z}$ is not.
Appendix B Proofs
B.1 Proof of Claim 1
Proof.
when $k>0$ is some fixed integer, we know that the scores of all $k$ runs are $\leq g(Y)$, which has the probability $\operatorname{Pr}(E_{\leq y})^{k}$. As $y$ occurs, we have probability $\operatorname{Pr}(E_{\leq y})^{k}-\operatorname{Pr}(E_{<y})^{k}$ seeing $y$ as the output of Algorithm 2. When $k$ follows some general distribution $\xi$, the resultant distribution is a mixture, which is Claim 1. ∎
B.2 Proof of Theorem 2
Proof.
W.o.l.g., we define the alphabets of distribution $P$ and $P^{\prime}$ as $\{a,b,c,d,e,f\}$, with some abuse of notation, we denote
$P(a)=p_{a},P(b)=p_{b},\cdots,P(f)=p_{f}$
$P^{\prime}(a)=p^{\prime}_{a},P^{\prime}(b)=p^{\prime}_{b},\cdots,P^{\prime}(f)=p^{\prime}_{f}$
as their probabilities. Suppose we have a non-one-to-one mapping score evaluator $\hat{g}$ such that:
$\hat{g}(a)=\hat{g}(c)=\hat{g}(e)<\hat{g}(b)=\hat{g}(d)<\hat{g}(f).$
We now assume a uniformly random selection among Alphabets that share the same score. For clearer presentation, we denote $\Lambda_{S}^{k}=(\sum_{i\in S}p_{i})^{k}$ and $\bar{\Lambda}_{S}^{k}=(\sum_{i\in S}p^{\prime}_{i})^{k}$. Then, the distribution of $F_{k,\hat{g}}$ will be
$F_{k,\hat{g}}(a)=F_{k,\hat{g}}(c)=F_{k,\hat{g}}(e)=\frac{1}{3}\Lambda_{\{a,c,e\}}^{k}$
$F_{k,\hat{g}}(b)=F_{k,\hat{g}}(d)=\frac{1}{2}(\Lambda_{\{a,c,e,b,d\}}^{k}-\Lambda_{\{a,c,e\}}^{k})$
$F_{k,\hat{g}}(f)=1-\Lambda_{\{a,c,e,b,d\}}^{k}$
This is because
$F_{k,\hat{g}}(i\in\{a,c,e\})=\Lambda_{\{a,c,e\}}^{k}$
$F_{k,\hat{g}}(i\in\{b,d\})=\Lambda_{\{a,c,e,b,d\}}^{k}-\Lambda_{\{a,c,e\}}^{k}$
and a uniformly random selection among $\{a,c,e\}$ means that the probability mass $F_{k,\hat{g}}(i\in\{a,c,e\})$ is distributed uniformly to each. Similarly, the distribution of $F^{\prime}_{k,\hat{g}}$ corresponding to $P^{\prime}$ has the same form (just replace $\Lambda_{S}^{k}$ by $\bar{\Lambda}_{S}^{k}$). We now construct a one-to-one mapping score function $g^{*}$ as follows.
$g^{*}(a)<g^{*}(c)<g^{*}(e)<g^{*}(b)<g^{*}(d)<g^{*}(f)$
The key point here is to enforce a strict total order for alphabets that have the same score. Then, the distribution of $F_{k,g^{*}}$ is
$F_{k,g^{*}}(a)=\Lambda_{\{a\}}^{k},F_{k,g^{*}}(c)=\Lambda_{\{a,c\}}^{k}-\Lambda_{\{a\}}^{k}$
$F_{k,g^{*}}(e)=\Lambda_{\{a,c,e\}}^{k}-\Lambda_{\{a,c\}}^{k},F_{k,g^{*}}(b)=\Lambda_{\{a,c,e,b\}}^{k}-\Lambda_{\{a,c,e\}}^{k}$
$F_{k,g^{*}}(d)=\Lambda_{\{a,c,e,b,d\}}^{k}-\Lambda_{\{a,c,e,b\}}^{k},F_{k,g^{*}}(f)=1-\Lambda_{\{a,c,e,b,d\}}^{k}$
Similarly, the distribution of $F^{\prime}_{k,g^{*}}$ corresponding to $P^{\prime}$ has the same form (just replace $\Lambda_{S}^{k}$ by $\bar{\Lambda}_{S}^{k}$). We now compute the RDP quantity $\mathcal{D}_{\alpha}(F_{k,\hat{g}}||F^{\prime}_{k,\hat{g}})$ and $\mathcal{D}_{\alpha}(F_{k,g^{*}}||F^{\prime}_{k,g^{*}})$. We aim to show that the RDP value under non-one-to-one mapping $\hat{g}$ is smaller than that under its one-to-one mapping counterpart. We group the sub-terms of RDP calculation. Let
$\hat{T}_{\{a,c,e\}}=\sum_{i\in\{a,c,e\}}\left(\frac{F_{k,\hat{g}}(i)}{F^{\prime}_{k,\hat{g}}(i)}\right)^{\alpha}F^{\prime}_{k,\hat{g}}(i)$
$\hat{T}_{\{b,d\}}=\sum_{i\in\{b,d\}}\left(\frac{F_{k,\hat{g}}(i)}{F^{\prime}_{k,\hat{g}}(i)}\right)^{\alpha}F^{\prime}_{k,\hat{g}}(i)$
$\hat{T}_{\{f\}}=\left(\frac{F_{k,\hat{g}}(f)}{F^{\prime}_{k,\hat{g}}(f)}\right)^{\alpha}F^{\prime}_{k,\hat{g}}(f)$
We compute the $T^{*}_{\{a,c,e\}},T^{*}_{\{b,d\}},T^{*}_{\{f\}}$ counterparts in the same fashion (just replace $\hat{g}$ by $g^{*}$). And we will compare $T_{\{a,c,e\}}$ and $T^{\prime}_{\{a,c,e\}}$. By letting
$x=\Lambda_{\{a\}}^{k}\text{ \quad\quad\quad\quad\quad\quad\quad}$ $x^{\prime}=\bar{\Lambda}_{\{a\}}^{k}$
$y=\Lambda_{\{a,c\}}^{k}-\Lambda_{\{a\}}^{k}\text{ \quad\quad\quad}$ $y^{\prime}=\bar{\Lambda}_{\{a,c\}}^{k}-\bar{\Lambda}_{\{a\}}^{k}$
$z=\Lambda_{\{a,c,e\}}^{k}-\Lambda_{\{a,c\}}^{k}\text{ \quad\quad}$ $z^{\prime}=\bar{\Lambda}_{\{a,c,e\}}^{k}-\bar{\Lambda}_{\{a,c\}}^{k}$
then, it is easy to see that
$\hat{T}_{\{a,c,e\}}=$ $(\frac{\Lambda_{\{a,c,e\}}^{k}}{\bar{\Lambda}_{\{a,c,e\}}^{k}})^{\alpha}\bar{\Lambda}_{\{a,c,e\}}^{k}$ (28)
$=$ $(\frac{x+y+z}{x^{\prime}+y^{\prime}+z^{\prime}})^{\alpha}(x^{\prime}+y^{\prime}+z^{\prime})$
$\leq$ $(\frac{x}{x^{\prime}})^{\alpha}x^{\prime}+(\frac{y}{y^{\prime}})^{\alpha}y^{\prime}+(\frac{z}{z^{\prime}})^{\alpha}z^{\prime}$
$=$ $T^{*}_{a,c,e}$
holds by Jensen’s inequality and the fact that function $h(x)=x^{\alpha}$ is convex for $\alpha>1$, i.e.,
$(\frac{x+y+z}{x^{\prime}+y^{\prime}+z^{\prime}})^{\alpha}\leq\frac{(\frac{x}{x^{\prime}})^{\alpha}x^{\prime}+(\frac{y}{y^{\prime}})^{\alpha}y^{\prime}+(\frac{z}{z^{\prime}})^{\alpha}z^{\prime}}{x^{\prime}+y^{\prime}+z^{\prime}}$
For the same reason, it can also be easily checked that $\hat{T}_{\{b,d\}}\leq T^{*}_{\{b,d\}}$ and $\hat{T}_{\{f\}}\leq T^{*}_{\{f\}}$ also hold. Because
$\mathcal{D}_{\alpha}(F_{k,\hat{g}}||F^{\prime}_{k,\hat{g}})=\frac{1}{\alpha-1}\ln(\hat{T}_{\{a,c,e\}}+\hat{T}_{\{b,d\}}+\hat{T}_{\{f\}})$
$\mathcal{D}_{\alpha}(F_{k,g^{*}}||F^{\prime}_{k,g^{*}})=\frac{1}{\alpha-1}\ln(T^{*}_{\{a,c,e\}}+T^{*}_{\{b,d\}}+T^{*}_{\{f\}}),$
we have
$\mathcal{D}_{\alpha}(F_{k,\hat{g}}||F^{\prime}_{k,\hat{g}})\leq\mathcal{D}_{\alpha}(F_{k,g^{*}}||F^{\prime}_{k,g^{*}}).$
Note the first equality of Equation (28) always holds no matter whether selection among alphabets sharing the same score is uniform or weighted; the alphabet and the order we choose is also arbitrary, which means that the result holds in general.
Remark. Following the same reasoning, when $k$ is now a random variable instead of a fixed number, we also have the result, as shown in the above theorem. Because we can modify each probability term to be the probability of the mixture counterpart, and the proof follows trivially. Specifically, for each probability $p=f(k)$ shows up, modify it to be $p=\sum_{i=1}^{\infty}\operatorname{Pr}(i)f(i)$ where $\operatorname{Pr}(i),i=\{1,2,\cdots,\infty\}$ is the p.m.f. of distribution $\xi$. ∎
B.3 Proof of Theorem 3
Proof.
As we care about how much $(\varepsilon_{\mathcal{H}},\delta_{\mathcal{H}})$-DP $\mathcal{H}$ satisfies given some $\delta_{\mathcal{H}}$, it is useful to introduce a technical lemma related to such form of DP.
Lemma 1 ([46] Propostion 7).
[…]
$\int_{\varepsilon}^{\infty}\mathrm{e}^{\varepsilon-z}\cdot\operatorname{Pr}_{o\sim\mathcal{M}(X)}[L_{1}>z]\mathrm{d}z\leq\delta$
holds for all adjacent $X,X^{\prime}$.
Our goal is clear, i.e., we need to meet the following equation:
$\int_{\varepsilon_{\mathcal{H}}}^{\infty}\mathrm{e}^{\varepsilon_{\mathcal{H}}-z}\cdot\operatorname{Pr}_{o\sim Q}[\log\frac{Q(o)}{Q^{\prime}(o)}>z]\mathrm{d}z\leq\delta_{\mathcal{H}}$ (29)
Then $\mathcal{H}$ would be $(\varepsilon_{\mathcal{H}},\delta_{\mathcal{H}})$-DP.
Let the left-hand side of Equation (29) to be $t_{\varepsilon_{\mathcal{H}}}$ and note that $\frac{Q(o)}{Q^{\prime}(o)}=\frac{P(o)\omega_{\xi}(F(o))}{P^{\prime}(o)\omega_{\xi}(F^{\prime}(o))}$, define event $E_{z}=\{o:\log\frac{P(o)\omega_{\xi}(F(o))}{P^{\prime}(o)\omega_{\xi}(F^{\prime}(o))}>z\}$ then
$t_{\varepsilon_{\mathcal{H}}}=$ $\int_{\varepsilon_{\mathcal{H}}}^{\infty}\mathrm{e}^{\varepsilon_{\mathcal{H}}-z}\cdot\int\limits_{E_{z}}Q(o)\mathrm{d}o\mathrm{d}z$ (30)
$\leq$ $\omega_{\xi}(1)\int_{\varepsilon_{\mathcal{H}}}^{\infty}\mathrm{e}^{\varepsilon_{\mathcal{H}}-z}\cdot\int\limits_{E_{z}}P(o)\mathrm{d}o\mathrm{d}z$
The inequality is due to $\omega_{\xi}:[0,1]\rightarrow\mathbb{R}$ is increasing.
Let us investigate the hypothesis testing problem $P$ V.S. $P^{\prime}$, i.e., deciding $X$ or $X^{\prime}$ was used based on the score of a single run of the DP-SGD. The score is post-processing [17, Lemma 1]) of the trained model, so the $(\mathrm{FP},\mathrm{FN})$ pair for distinguishing $P$ from $P^{\prime}$ is governed by $f$.
For some real number $o\in\mathbb{R}$, define $A=\{u:u\leq o\}$, and a decision rule $\mathcal{R}$ that accepts $P$ when the score falls into $A$. Then, $\mathrm{FP}_{\mathcal{R}}=1-F(o)$ and $\mathrm{FN}_{\mathcal{R}}=F^{\prime}(o)$. And we must have $F^{\prime}(o)\geq f(1-F(o))$ as governed by the trade-off function. This leads to an upper bound (note that $\omega_{\xi}$ is increasing)
$\frac{\omega_{\xi}(F(o))}{\omega_{\xi}(F^{\prime}(o))}\leq\max_{a\in[0,1]}{\frac{\omega_{\xi}(1-a)}{\omega_{\xi}(f(a))}}=M$ (31)
Now, let $\hat{E}_{z}=\{o:\log\frac{P(o)}{P^{\prime}(o)}M>z\}$, it is easy to see that $E_{z}\subseteq\hat{E}_{z}$. Hence, we have
$t_{\varepsilon_{\mathcal{H}}}\leq$ $\omega_{\xi}(1)\int_{\varepsilon_{\mathcal{H}}}^{\infty}\mathrm{e}^{\varepsilon_{\mathcal{H}}-z}\cdot\int\limits_{\hat{E}_{z}}P(o)\mathrm{d}o\mathrm{d}z$
$=$ $\omega_{\xi}(1)\int_{\varepsilon_{\mathcal{H}}}^{\infty}\mathrm{e}^{\varepsilon_{\mathcal{H}}-z}\cdot\operatorname{Pr}_{o\sim P}[\log\frac{P(o)}{P^{\prime}(o)}>z-\log M]\mathrm{d}z$
$\leq$ $\omega_{\xi}(1)\int_{\varepsilon_{\mathcal{H}}-\log M}^{\infty}\mathrm{e}^{\varepsilon_{\mathcal{H}}-z}\cdot\operatorname{Pr}_{o\sim P}[\log\frac{P(o)}{P^{\prime}(o)}>z-\log M]\mathrm{d}z$
$=$ $\omega_{\xi}(1)\int_{\varepsilon_{\mathcal{H}}-\log M}^{\infty}\mathrm{e}^{\varepsilon_{\mathcal{H}}-\log M-s}\cdot\operatorname{Pr}_{o\sim P}[\log\frac{P(o)}{P^{\prime}(o)}>s]\mathrm{d}s$
Letting $s=z-\log M$, we have the last equality. Note that the score is differentially private, as it is post-processing of the base algorithm. Hence, we can compute a $(\varepsilon_{\mathcal{H}}-\log M,\delta)$-DP guarantee for the score. Applying Lemma 1, we have
$t_{\varepsilon_{\mathcal{H}}}\leq\omega_{\xi}(1)\delta.$
Setting $\delta_{\mathcal{H}}=\omega_{\xi}(1)\delta$, we derive $\delta$. By inputting trade-off function $f$ for the base algorithm and $\delta$ to Algorithm 3, we derive the value of $\epsilon_{\mathcal{H}}-\log{M}$, which give us Theorem 3. As we assume nothing on the adjacent dataset $X,X^{\prime}$, Theorem 3 holds for all $X,X^{\prime}$ pair.